### v1

# Robust and Computation-Aware Gaussian Processes

Abstract
Gaussian processes (GPs) are widely used for regression and optimization tasks such as Bayesian optimization (BO) due to their expressiveness and principled uncertainty estimates. However, in settings with large datasets corrupted by outliers, standard GPs and their sparse approximations struggle with computational tractability and robustness. We introduce Robust Computation-aware Gaussian Process (RCaGP), a novel GP model that jointly addresses these challenges by combining a principled treatment of approximation-induced uncertainty with robust generalized Bayesian updating. The key insight is that robustness and approximation-awareness are not orthogonal but intertwined: approximations can exacerbate the impact of outliers, and mitigating one without the other is insufficient. Unlike previous work that focuses narrowly on either robustness or approximation quality, RCaGP combines both in a principled and scalable framework, thus effectively managing both outliers and computational uncertainties introduced by approximations such as low-rank matrix multiplications. Our model ensures more conservative and reliable uncertainty estimates, a property we rigorously demonstrate. Additionally, we establish a robustness property and show that the mean function is key to preserving it, motivating a tailored model selection scheme for robust mean functions. Empirical results confirm that solving these challenges jointly leads to superior performance across both clean and outlier-contaminated settings, both on regression and high-throughput Bayesian optimization benchmarks.
1 Introduction
Gaussian Processes (GPs) are a foundational tool in probabilistic machine learning, offering non-parametric modeling with principled uncertainty estimates [35]. Their use spans diverse domains such as time series forecasting, spatial modeling, and regression on structured scientific data. One particularly impactful area where GPs have become integral is Bayesian Optimization (BO), a sample-efficient framework for optimizing expensive black-box functions, where the surrogate model’s predictive accuracy and uncertainty calibration directly influence decision-making [10]. Many scientific and industrial tasks are high-throughput optimization problems in which numerous experiments or simulations are performed, possibly in parallel. These can be effectively tackled by BO and include material design [14], drug discovery with extensive chemical space search [17], and fine-tuning gene expression in synthetic biology [5].
Yet, two major challenges limit the reliability of GPs in both regression and BO settings: scalability to large datasets and robustness to corrupted or outlier-contaminated observations. Sparse Variational Gaussian Processes (SVGPs, [16, 40, 41]) address scalability via inducing point approximations, yielding a low-rank approximation that allows large-scale regression and high-throughput BO.
Figure 1: Overview of our proposed RCaGP against concurrent baselines on a 1D example. (Left) SVGP fails to fit observed data contaminated by outliers, whereas RCaGP successfully fits high-density data regions while preserving a higher variance due to the presence of outliers. (Middle-left) While enhancing robustness compared to SVGP, RCSVGP deviates more significantly from the true function than RCaGP. (Middle-right) Even if CaGP displays increased posterior variance near outliers, RCaGP provides superior posterior mean prediction. (Right) As a result, the acquisition function landscape for RCaGP better prioritizes the true global optima.
Despite their computational efficiency, significant challenges hinder the adoption of SVGPs for high-throughput BO tasks, due to the impact of surrogate posterior quality on data acquisition. For starters, SVGP is not robust against outliers: upon contamination of the observed data, the posterior mean can deviate significantly from the true latent function. Such deviation will misguide most acquisition strategies, resulting in suboptimal performance. Outliers are an almost unavoidable aspect of many real-world applications and datasets, specifically in high-throughput settings. They can arise from various factors, such as faulty measurements, malfunctioning sensors, or genetic anomalies. Next, SVGPs suffer from overconfident predictions, a well-known pathology caused by the error incurred by the posterior approximation. Overconfident predictions occur in regions where the inducing points are far from observed data. Accurate uncertainty quantification is essential for sequentially identifying the optimal data points in BO [44].
Contributions. To address these challenges, we propose a novel approach that combines robust-conjugate SVGP [2] with computation-aware GP [46]. This method, referred to as the Robust Computation-Aware GP (RCaGP), jointly tackles outlier robustness and approximation-induced uncertainty—two challenges that are deeply intertwined in sparse GP models. Computational approximations such as low-rank matrix factorizations can exacerbate the effects of outliers, and mitigating one without addressing the other leads to suboptimal performance. RCaGP can be used as a surrogate in high-throughput BO, and its design is motivated by the need to address these two problems in tandem. This can be seen in Figure 1. In summary, this work
  1. 1.
    (Methodological) introduces a novel GP model (RCaGP) that jointly addresses approximation-awareness and outlier robustness through a principled probabilistic framework—going beyond previous work that tackles these issues in isolation.
  2. 2.
    (Theoretical) establishes RCaGP’s reliability through conservative uncertainty estimates and a formal robustness property. Our analysis reveals the central role of the mean function in preserving robustness, motivating a tailored model selection scheme.
  3. 3.
    (Empirical) demonstrates that jointly solving these two problems yields superior performance: RCaGP consistently outperforms existing state-of-the-art methods such as CaGP and RCSVGP on UCI regression tasks and BO settings, with and without outliers.
2 Background
Robust conjugate GP. The Gaussian noise assumption in GPs often leads to misspecification, e.g., due to outliers. To address this problem, recent work proposed robust conjugate GP (RCGP), which employs a robust loss function through generalized Bayesian inference [2]. Consider a noisy dataset ${\mathcal{D}}_{n}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}=({\bf{X}},{\bf{y}})$. We specify a GP prior for the finite function evaluations ${\bf{f}}=[f({\bf{x}}_{1}),\dots,f({\bf{x}}_{n})]$: $p({\bf{f}})={\mathcal{N}}({\bf{f}};\!{\bf{m}},\!{\bf{K}})$, with ${\bf{m}}=[m({\bf{x}}_{1}),\dots,m({\bf{x}}_{n})]^{\top}$ and ${\bf{K}}=[k({\bf{x}}_{j},\!{\bf{x}}_{l})]_{1\leq j,l\leq n}$. Both $m:{\mathcal{X}}\rightarrow$ and $k:{\mathcal{X}}\times{\mathcal{X}}\rightarrow$ denote the mean and kernel functions, encoding prior beliefs about the function’s structure and smoothness [35]. Conditioning on observations, the posterior predictive distribution is Gaussian with mean $\mu_{\ast}({\bf{x}})$ and variance $k_{\ast}({\bf{x}},\!{\bf{x}})$.
$\mu_{\ast}({\bf{x}})=m({\bf{x}})+{\bf{k}}_{\bf{x}}^{\top}({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}}),$ (1)
$k_{\ast}({\bf{x}},{\bf{x}})=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}{\bf{k}}_{\bf{x}},$ (2)
where ${\bf{k}}_{\bf{x}}=[k({\bf{x}},{\bf{x}}_{1}),\dots k({\bf{x}},{\bf{x}}_{n})]^{\top}\in^{n}$. The key difference with vanilla GPs is the introduction of a weight function $w$, used by RCGP to down-weight the influence of potential outliers. Concretely, RCGP induces a corrected mean and posterior variance by replacing $({\bf{y}}-{\bf{m}})$ with the shrinkage term $({\bf{y}}-{\bf{m}}_{\bf{w}})$, with ${\bf{m}}_{\bf{w}}={\bf{m}}+{\sigma_{\mathrm{noise}}^{2}}\nabla_{y}\log({\bf{w}}^{2})$ and the identity matrix ${\bf{I}}$ with a diagonal matrix ${\bf{J}}_{\bf{w}}=\mathrm{diag}\left(\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}{\bf{w}}^{-2}\right)$, for ${\bf{w}}=[w({\bf{x}}_{1},y_{1}),\dots,w({\bf{x}}_{n},y_{n})]^{\top}$.
Nevertheless, RCGP inherits the cubic inference complexity of vanilla GPs, prohibiting its usage on large-scale datasets like those encountered in high-throughput BO. To remedy this issue, RCSVGP were introduced, enabling outlier-robust regression on large-scale data [2]. This model introduces inducing points ${\bf{u}}=\{{\bf{u}}_{l}\}_{l=1}^{r}$ for $r\ll n$, with inducing locations ${\bf{z}}=\{{\bf{z}}_{l}\}_{l=1}^{r}$ such that ${\bf{u}}_{l}=f({\bf{z}}_{l})$. Given ${\bf{u}}$, RCSVGP constructs an approximate posterior $q({\bf{u}})={\mathcal{N}}({\boldsymbol{\mu}}_{\bf{u}},{\boldsymbol{\Sigma}}_{\bf{u}})$ with variational parameters ${\boldsymbol{\mu}}_{\bf{u}}$ and ${\boldsymbol{\Sigma}}_{\bf{u}}$, obtained by maximizing the evidence lower bound (ELBO) criterion:
$\ell_{\mathrm{ELBO}}^{\mathrm{RCSVGP}}={\mathbb{E}}_{q({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]-{\mathrm{KL}}[q({\bf{u}})|p({\bf{u}})],$ (3)
where $\Psi^{w}({\bf{y}},{\bf{f}})$ denotes the robust loss function and $q({\bf{f}})=p({\bf{f}}|{\bf{u}})q({\bf{u}})$ is the approximate posterior over the function values conditioned on ${\bf{u}}$.
While RCSVGP reduces the time complexity to ${\mathcal{O}}(nr^{2}),$ like other approximate GP methods, it also exhibits overconfidence in regions where inducing points are far from the observations. This overconfidence can harm the exploration/exploitation balance in BO, leading to suboptimal results.
Computation-aware GP. [45, 46] proposed computation-aware GP (CaGP) to address the SVGP overconfidence. CaGP mitigates this issue by ensuring its posterior variance exceeds the vanilla GP variance. This corrected variance is obtained by adding computational uncertainty, associated with the approximation error in the representer weights ${\bf{v}}_{\ast}=({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})^{-1}({\bf{y}}-{\bf{m}})$. To obtain computational uncertainty, CaGP places a prior $p({\bf{v}}_{\ast})={\mathcal{N}}({\bf{v}}_{0},{\boldsymbol{\Sigma}}_{0})$ and updates its posterior distribution $\mathcal{N}({\bf{v}}_{i},{\boldsymbol{\Sigma}}_{i})$ using a probabilistic linear solver based on linear Gaussian identities:
${\bf{v}}_{i}={\bf{C}}_{i}\,({\bf{y}}-{\bf{m}}),$ (4)
${\boldsymbol{\Sigma}}_{i}={\bf{K}}^{-1}-{\bf{C}}_{i},$ (5)
where ${\bf{C}}_{i}={\bf{S}}_{i}({\bf{S}}_{i}^{\top}\hat{{\bf{K}}}^{-1}{\bf{S}}_{i}){\bf{S}}_{i}\approx{\bf{K}}^{-1}$ is a low-rank approximation and ${\bf{S}}_{i}\in^{n\times i}$ is the action matrix associated with the chosen approximation technique. In the case of inducing points, ${\bf{S}}_{i}=k({\bf{X}},{\bf{z}}_{i})$. Given a new data point ${\bf{x}}$, the predictive distribution of $f({\bf{x}})$ is Gaussian, characterized by the mean $\mu_{i}({\bf{x}})$ and variance $k_{i}({\bf{x}},{\bf{x}})$, given by:
$\mu_{i}({\bf{x}})=m({\bf{x}})+{\bf{k}}_{\bf{x}}^{\top}{\bf{v}}_{i},$ (6)
$k_{i}({\bf{x}},{\bf{x}})=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}\hat{{\bf{K}}}^{-1}{\bf{k}}_{\bf{x}}+{\bf{k}}_{\bf{x}}^{\top}{\boldsymbol{\Sigma}}_{i}{\bf{k}}_{\bf{x}}=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}{\bf{C}}_{i}{\bf{k}}_{\bf{x}},$ (7)
where $\hat{{\bf{K}}}={\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{I}}$ and $\sigma_{i}({\bf{x}})={\bf{k}}_{\bf{x}}^{\top}{\boldsymbol{\Sigma}}_{i}{\bf{k}}_{\bf{x}}$ denote the computational uncertainty. The posterior can be obtained in ${\mathcal{O}}(n^{2}i)$ time complexity. As $i\rightarrow n$, CaGP recovers both the predictive mean and variance of the exact GP.
Bayesian optimization. BO is a sequential algorithm designed to solve the problem ${\mathrm{argmax}}_{{\bf{x}}\in{\mathcal{X}}}f({\bf{x}})$ for ${\mathcal{X}}$ a compact domain, and $f$ a black-box function, meaning only noisy evaluations are accessible [10]. BO involves a probabilistic surrogate $p(f|{\mathcal{D}})$ and an acquisition function (AF) $\alpha({\bf{x}}):=\int u({\bf{x}};f)\,p(f|{\mathcal{D}})\,\mathrm{d}f$, where $u$ measures the usefulness of evaluating $f({\bf{x}})$. GPs are usually employed as the surrogate, and popular AFs include, for instance, expected improvement, knowledge gradient, and entropy search [10]. Importantly, these AFs can be maximized in a batch setting, yielding $q$ candidates every iteration $t,{\bf{X}}_{t}^{q}={\mathrm{argmax}}_{{\bf{X}}\in{\mathcal{X}}^{q}}\alpha({\bf{X}})$ [48].
3 Robust and Computation-Aware Gaussian Processes
To tackle the intertwined challenges of outliers and approximation-induced overconfidence, we introduce RCaGP—a principled integration of robust-conjugate GPs with computation-aware approximations. These issues must be addressed jointly: approximations can amplify outlier effects, while robustness alone does not correct the biases they introduce. RCaGP unifies both aspects to improve predictive reliability and decision-making. Section 3.1 details its composite inference scheme, Section 3.2 derives a tailored optimization criterion, and Section 3.3 extends the framework to BO for joint model and query selection.
3.1 Robust Computation-aware Inference
Inspired by CaGP, we apply a probabilistic treatment to the representer weights $\hat{{\bf{v}}}=({\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}\!-\!{\bf{m}}_{\bf{w}})$ of RCGP in Equation 1. This enables us to quantify uncertainty in approximating $\hat{{\bf{v}}}$—an aspect most approximate GPs ignore. RCaGP later incorporates this uncertainty into its predictive variance. We begin by placing a prior on $\hat{{\bf{v}}}$, i.e., $\hat{{\bf{v}}}\sim{\mathcal{N}}(\tilde{{\bf{v}}}_{0}\!=\!\mathbf{0},\,\tilde{{\boldsymbol{\Sigma}}}_{0}\!=\!\tilde{{\bf{K}}}^{-1})$, where $\tilde{{\bf{K}}}={\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}}$. We then update this belief via the linear Gaussian identity [45], yielding a Gaussian posterior with mean $\tilde{{\bf{v}}}_{i}$ and variance $\tilde{{\boldsymbol{\Sigma}}}_{i}$:
$\tilde{{\bf{v}}}_{i}=\tilde{{\bf{C}}}_{i}\,({\bf{y}}-{\bf{m}}_{\bf{w}}),$ (8)
$\tilde{{\boldsymbol{\Sigma}}}_{i}=\tilde{{\bf{K}}}^{-1}-\tilde{{\bf{C}}}_{i},$ (9)
where $\tilde{{\bf{C}}}_{i}\!=\!{\bf{S}}_{i}({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}$ and ${\bf{S}}_{i}\!\in\!^{n\times i}$ are low-rank approximations of $\tilde{{\bf{K}}}^{-1}$ and the actions, respectively. The updates can be done sequentially [45, Algorithm 1] or in batches [46, Algorithm S2].
Recent work generalized the action matrix ${\bf{S}}_{i}\in^{n\times i}$ by treating it as a variable that can be optimized during model selection [46]. Given a new data point ${\bf{x}}\in{\mathcal{X}}$, the predictive distribution of $f({\bf{x}})$ follows a normal distribution with mean $\hat{\mu}_{i}({\bf{x}})$ and variance $\hat{k}_{i}({\bf{x}},{\bf{x}})$:
$\hat{\mu}_{i}({\bf{x}})=m({\bf{x}})+{\bf{k}}_{\bf{x}}^{\top}\tilde{{\bf{v}}}_{i},$ (10)
$\hat{k}_{i}({\bf{x}},{\bf{x}})=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}\tilde{{\bf{K}}}_{i}{\bf{k}}_{\bf{x}}+{\bf{k}}_{\bf{x}}^{\top}\tilde{{\boldsymbol{\Sigma}}}_{i}{\bf{k}}_{\bf{x}}=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}\tilde{{\bf{C}}}_{i}{\bf{k}}_{\bf{x}}.$ (11)
In Equation 11, RCaGP incorporates the computational uncertainty ${\bf{k}}_{\bf{x}}^{\top}\,\tilde{{\boldsymbol{\Sigma}}}_{i}\,{\bf{k}}_{\bf{x}}$, corresponding to the variance of the representer weights $\tilde{{\boldsymbol{\Sigma}}}_{i}$, resulting in a combined uncertainty $\hat{k}_{i}({\bf{x}},{\bf{x}})$. This computational uncertainty guarantees that the predictive posterior variance of RCaGP is larger than RCGP, preventing RCaGP from overconfident predictions. Such behavior is referred to as a conservative uncertainty estimate. The computational complexity of RCaGP can be found in Appendix B.
The robustness of RCaGP hinges on the weight function $w$, which we choose to incorporate into the representer weights $\tilde{{\bf{v}}}$. To that end, we follow [2] and let:
$w({\bf{x}},y)=\beta\left(1+(y-m({\bf{x}}))^{2}/c^{2}\right)^{-1/2},$ (12)
for $\beta$ and $c>0$ the learning rate and soft threshold.
The weight function assigns smaller values to observations $y$ that deviate significantly from the mean prior $m({\bf{x}})$, treating them as potential outliers. Lower weights reduce the influence of such points during inference, leading to a more robust posterior. Additionally, the weight enters the noise term ${\bf{J}}_{\bf{w}}=\mathrm{diag}([{\sigma_{\mathrm{noise}}^{2}}/2{\bf{w}}^{-2}])$, effectively inflating the noise for outliers and further limiting their impact. $w({\bf{x}},y)$ depends critically on the mean prior: for an outlier $\hat{y}$, if $m(\hat{{\bf{x}}})$ is close to $\hat{y}$, the weight increases. However, choosing an informative mean prior is difficult without domain knowledge.
3.2 Model hyperparameters optimization
We use the evidence lower bound (ELBO) as a loss function to optimize the kernel hyperparameters ${\boldsymbol{\theta}}\!\in\!^{p}$. Following [46], this enables RCaGP to scale to large datasets while avoiding overconfidence. The variational family is defined using the RCaGP posterior $q_{i}({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})$, i.e., $\{q_{i}({\bf{f}})={\mathcal{N}}(\hat{\mu}_{i}({\bf{X}}),\hat{k}_{i}({\bf{X}},\!{\bf{X}}))|{\bf{S}}_{i}\!\in\!^{n\times i}\}$, parameterized by action ${\bf{S}}_{i}$. We replace the evidence term $\log p({\bf{y}}|{\boldsymbol{\theta}})$ with the robust loss $\log\Psi^{w}({\bf{y}},\!{\bf{f}})$ from [2]. We then formulate ELBO as
$\ell_{\mathrm{ELBO}}^{\mathrm{RCaGP}}={\mathbb{E}}_{q_{i}({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]-{\mathrm{KL}}[q_{i}({\bf{f}})\|p({\bf{f}})].$ (13)
This loss learns the hyperparameters as if maximizing ${\mathbb{E}}_{p({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ while minimizing the computational uncertainty associated with the approximation error. The derivation of the expected loss ${\mathbb{E}}_{q({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ follows [2], while the derivation of the KL term ${\mathrm{KL}}[q({\bf{f}})\|p({\bf{f}})]$ draws an analogy to CaGP. For the closed form and detailed derivation of ELBO, we refer the readers to Proposition C.1. We obtain the optimal hyperparameters ${\boldsymbol{\theta}}^{\ast}$ by minimizing the negative ELBO.
3.3 Joint model parameters and design selection strategy
To fully take advantage of RCaGP in BO, we leverage the expected lower bound utility (EULBO) framework introduced by [28]. At acquisition time, instead of optimizing the surrogate’s inducing points, variational parameters, and hyperparameters, only then to find a design maximizing the acquisition function, a global criterion is maximized in an end-to-end manner:
$\ell_{\mathrm{EULBO}}^{\mathrm{RCaGP}}=\ell_{\mathrm{ELBO}}^{\mathrm{RCaGP}}+{\mathbb{E}}_{q_{i}({\bf{f}})}[\log u({\bf{x}},f;{\mathcal{D}}_{t})],$ (14)
where $u:{\mathcal{X}}\times{\mathcal{F}}\rightarrow$ is the utility function. Here, we consider the expected improvement (EI) AF, reformulated as an expectation of the following utility:
$u_{\mathrm{EI}}({\bf{x}},f;{\mathcal{D}}_{t})=\mathrm{softplus}(f({\bf{x}})-y_{t}^{\ast}),$ (15)
where $y_{t}^{\ast}$ denotes the best evaluation observed so far and $\mathrm{softplus}:{\bf{x}}\mapsto\log(1+\exp({\bf{x}}))$. This reformulation replaces the commonly used ReLU function with a softplus function, guaranteeing the utility function remains strictly positive whenever $f({\bf{x}})\geq y^{\ast}$. Moreover, EULBO can be extended to support batch BO using Monte Carlo batch mode [4, 47]. Given a set of candidates ${\bf{X}}=\{{\bf{x}}_{1},\dots,{\bf{x}}_{q}\}$, the expected utility function corresponding to the $q-$improvement utility is:
${\mathbb{E}}[\log u_{\mathrm{EI}}({\bf{X}},f;{\mathcal{D}}_{t})]\approx\frac{1}{S}\sum_{s=1}^{S}\underset{s=1,\dots,q}{\max}\mathrm{softplus}(r_{s}),$ (16)
with $r_{s}=f({\bf{x}}_{s})+\epsilon_{s}-y_{t}^{\ast}$ and $\epsilon_{s}\sim{\mathcal{N}}(0,1)$. Finally, when considering RCaGP, the action matrices ${\bf{S}}_{i}$ are optimized, leading to the joint maximization problem
${\bf{x}}_{1}^{\ast},\dots,{\bf{x}}_{q}^{\ast},{\bf{S}}_{i}^{\ast},{\boldsymbol{\theta}}_{t}^{\ast}=\underset{{\bf{x}}_{1},\dots,{\bf{x}}_{q},{\bf{S}},{\boldsymbol{\theta}}}{{\mathrm{argmax}}}\;\ell_{\mathrm{EULBO}}^{\mathrm{RCaGP}}.$ (17)
For efficiency, we impose the sparse block structure on the action matrices ${\bf{S}}_{i}\in^{n\times i}$, following [46]. Specifically, we enforce each block to be column vector ${\bf{s}}_{j}\in^{k\times 1}$, with $k=n/i$ entries, so that the number of trainable parameters is $k\times i=n$ (training data size). It is worth noticing that unlike EULBO with SVGP, which maximizes variational parameters and inducing locations, our approach directly optimizes the action matrices ${\bf{S}}_{i}$.
Maximizing the ELBO independently of the posterior-expected utility function can result in suboptimal data acquisition decisions, as the ELBO is primarily designed to model observed data [27]. In contrast, EULBO considers how the surrogate performs when selecting the next query, effectively guiding the solution of ELBO optimization towards high utility regions [28]. Our RCaGP model further balances such aggressive behavior by incorporating the combined uncertainty, preventing overconfidence during candidate query selection. Moreover, RCaGP enhances robustness in scenarios where function evaluations occasionally produce outliers.
4 Theoretical Analysis
This section presents a theoretical analysis of RCaGP. Proposition 4.1 establishes its robustness to outliers, while Proposition 4.2 links its uncertainty estimates to worst-case error, ensuring conservative predictions, thus addressing the challenges outlined in Section 1.
We demonstrate the robustness of RCaGP through the posterior influence function (PIF), similarly to [2, Proposition 3.2]. The PIF is a criterion proving robustness to misspecification in observation error [11, 26]. For this purpose, we define the dataset ${\mathcal{D}}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}$ and the corresponding contaminated dataset ${\mathcal{D}}_{m}^{c}=({\mathcal{D}}\setminus({\bf{x}}_{m},y_{m}))\cup({\bf{x}}_{m},y_{m}^{c})$, indexed by $m\in\{1,\dots,n\}$. The impact of $y_{m}^{c}$ on inference is measured through PIF, expressed as Kullback-Leibler (KL) divergence between RCaGP’s contaminated posterior $p({\bf{f}}|{\mathcal{D}}_{m}^{c})$ and its uncontaminated counterpart $p({\bf{f}}|{\mathcal{D}})$:
$\mathrm{PIF}(y_{m}^{c},{\mathcal{D}})=\mathrm{KL}(p({\bf{f}}|{\mathcal{D}})\|p({\bf{f}}|{\mathcal{D}}_{m}^{c})).$ (18)
PIF views the KL-divergence as a function of $|y_{m}^{c}-y_{m}|$. In principle, we can consider any divergence that is not uniformly bounded. Here, we choose KL-divergence since it provides a closed-form expression for multivariate Gaussian. A posterior is robust if $\sup_{y\in\mathcal{Y}}|\mathrm{PIF}(y_{m}^{c},{\mathcal{D}})|<\infty$ [2]. We then show that RCaGP’s PIF is bounded through the following proposition:
Proposition 4.1.
Let ${\bf{f}}\sim{\mathcal{GP}}(m,k)$ denote the RCaGP prior, and let $i\in{0,\dots,n}$ represent the number of actions in RCaGP. Define constants $C_{1}^{\prime}$ and $C_{2}^{\prime}$, which are independent of $y_{m}^{c}$. For any $i$ and assuming $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$, the PIF of RCaGP is given by:
$\mathrm{PIF}_{\mathrm{RCaGP}}(y_{m}^{c},{\mathcal{D}})=C_{1}^{\prime}(w({\bf{x}}_{m},y_{m}^{c})^{2}y_{m}^{c})^{2}+C_{2}^{\prime}.$ (19)
Thus, if $\sup_{{\bf{x}},y}y\,w({\bf{x}},y)^{2}<\infty$, then RCaGP is robust since $\sup_{y_{m}^{c}}|\mathrm{PIF}_{\mathrm{RCaGP}}(y_{m}^{c},{\mathcal{D}})|<\infty$.
The constraint $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$ ensures no observation has infinite weight, and $\sup_{{\bf{x}},y}y,w^{2}({\bf{x}},y)<\infty$ guarantees that $w$ down-weights observations at least at rate $1/y$. The proof sketch shows that each term in Equation 18 is bounded. We leverage the pseudo-inverse property of positive semi-definite matrices and matrix norm bounds on the low-rank approximation to relate RCGP and CaGP. The full proof is in Appendix D. Crucially, the PIF of CaGP is unbounded, indicating a lack of robustness to outliers (see Section F.1).
Next, we show that RCaGP’s uncertainty captures the worst-case error over all latent functions, paralleling [45, Theorem 2]. We assume the difference between the latent function $g$ and the shrinkage term $m_{w}$ lies in an RKHS. Under this assumption, we establish the link between RCaGP’s uncertainty estimates and the worst-case error through the following proposition:
Proposition 4.2.
Let $\hat{k}_{i}(\cdot,\cdot)=k_{\ast}(\cdot,\cdot)+\hat{\sigma}_{i}(\cdot,\cdot)$ be the combined uncertainty of RCaGP with zero-mean prior $m$. Then, for any new ${\bf{x}}\in\mathcal{X}$ we have that
$\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}(h({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}}))^{2}=\hat{k}_{i}({\bf{x}},{\bf{x}})+{\sigma_{\mathrm{noise}}^{2}}$ (20)
$\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}(\mu_{\ast}^{g}({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}}))^{2}=\hat{\sigma}_{i}({\bf{x}},{\bf{x}})$ (21)
where $\mu_{\ast}^{g}(\cdot)=k(\cdot,{\bf{X}})\tilde{{\bf{K}}}^{-1}h({\bf{X}})$ is the RCGP’s posterior and $\hat{\mu}_{i}^{g}(\cdot)=k(\cdot,{\bf{X}})\tilde{{\bf{C}}}_{i}(g({\bf{X}})-m_{w}({\bf{X}}))$ RCaGP’s posterior mean for a function $h\in{\mathcal{H}}_{k^{w}}=g({\bf{x}})-m_{w}({\bf{x}})$ with a latent function $g:{\mathcal{X}}\rightarrow$ and the shrinkage function $m_{w}:{\mathcal{X}}\times{\mathcal{Y}}\rightarrow$.
The first equation in Proposition 4.2 shows that RCaGP’s uncertainty $\hat{k}_{i}(.,.)$ captures the worst-case error between the latent function $g$ and the posterior mean $\hat{\mu}_{i}(.)$. The second shows that the computational uncertainty $\hat{\sigma}_{i}^{2}$ captures the worst-case error between the RCGP and RCaGP posterior means, induced by approximation actions such as inducing points. This added uncertainty prevents overconfidence when modeling the latent function. The proof follows [45], leveraging [Proposition 3.9] of [20] and the RKHS associated with RCGP. Full details are in Appendix E. Lastly, the convergence of RCaGP in terms of mean function in RKHS norm is established in Section F.2.
5 Expert-guided robust mean prior definition
Our proposed RCaGP enforces outlier robustness through the weight function $w$, which—as discussed in Section 3.1 and Proposition 4.2—critically depends on the prior mean $m$. Motivated by these insights, we propose defining a robust and informative mean prior using domain expert feedback.
We assume the expert can identify a subset of outliers in ${\mathcal{D}}_{n}$ and provide corrections. Instead of discarding these outliers, we use them to inform the prior: removing them would increase evaluation cost, while RCaGP can still benefit from them. Since expert corrections are imperfect, we model them probabilistically and define the mean via the inferred posterior.
Let $\mathbf{o}=\{o_{j}\}_{j=1}^{n}$ be binary outlier labels for dataset ${\mathcal{D}}_{n}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}$, where $o_{j}=1$ indicates an outlier. Each $o_{j}$ follows a Bernoulli likelihood $p(o_{j}|\delta_{j})$ with latent probability $\delta_{j}\sim\mathcal{B}(\alpha_{j},\beta_{j})$. We infer $p(\delta_{j}|o_{j})$ via Bayes’ rule. Next, let $\bar{\mathbf{o}}=\{\bar{o}_{o}\}_{o=1}^{\hat{o}}$ and $\hat{{\bf{y}}}=\{\hat{y}_{o}\}_{o=1}^{\hat{o}}$ denote the identified outliers, and $\bar{{\bf{y}}}=\{\bar{y}_{o}\}_{o=1}^{\hat{o}}$ the expert-provided corrections. Each correction $\bar{y}_{o}$ is drawn from ${\mathcal{N}}(\bar{\mu}_{o},{\sigma_{\mathrm{corr.}}^{2}})$ with latent $\bar{\mu}_{o}\sim{\mathcal{N}}(\mu_{o},\tau_{o})$. Posterior inference yields $p(\bar{\mu}_{o}|\bar{y}_{o})$. We define the mean prior by combining the expected labels and corrected values using a kernel-weighted average:
$m({\bf{x}})=\frac{1}{\hat{n}}\sum_{\hat{j}=1}^{\hat{n}}{\mathbb{E}}_{p(\delta_{\hat{j}}|\bar{o}_{\hat{j}})}[\delta_{\hat{j}}]{\mathbb{E}}_{p(\bar{\mu}_{\hat{j}}|\bar{y}_{\hat{j}})}[\bar{\mu}_{\hat{j}}]l({\bf{x}},{\bf{x}}_{j}),$ (22)
This corresponds to averaging over the $\hat{n}$ nearest identified outliers, weighted by a kernel $l$ and the expected confidence in each correction. In practice, we average over all identified outliers and use only posterior expectations, prioritizing computational efficiency over full Bayesian inference. The full generative process and toy illustration appear in Figure 2 (left and middle). Additional inference details are provided in Appendix G.
Figure 2: Expert-guided robust mean prior. (Left) Graphical model of expert’s feedback generative process for identified outliers. (Middle) User graphical model for expert outlier corrections. (Right) Toy example: with expert corrections, the expert-guided prior better captures the true function than a constant mean prior under contamination.
6 Related Work
Outlier-robust Gaussian processes fall into three main categories. The first replaces Gaussian noise with heavy-tailed alternatives—e.g., Student-$t$[19, 34], Huber[1], Laplace [22], data-dependent [12], or mixture models [7, 24]. These approaches break conjugacy and require approximate inference. The second class removes suspected outliers before applying standard inference [23, 31, 3], but this is often unreliable and computationally expensive in high dimensions. The third approach, RCGP [2], builds on generalized Bayesian inference [6, 18, 21] to down-weight outliers without discarding data. We adopt RCGP as the foundation for RCaGP due to its strong performance in regression with outliers, its preserved conjugacy, and greater computational efficiency—enabling more accurate posterior inference and better predictive performance.
SVGPs are the standard approach for scaling Gaussian processes to large-scale Bayesian optimization [13, 37, 43, 42]. [30] improved inducing point selection by optimizing their weighted Gram matrix via a determinantal point process. [28] formulated acquisition as expected utility, enabling joint optimization of queries, variational parameters, hyperparameters, and inducing points in an end-to-end fashion. [25] proposed an online conditioning strategy that updates SVGPs with new data without reoptimizing variational parameters, enabling look-ahead acquisition in BO.
Computation-aware Gaussian processes are a novel class of GPs that accounts for computational uncertainty, due to approximations in matrix-vector multiplications [45]. These methods fall under the domain of probabilistic numerics, which quantify additional uncertainties arising from approximation errors [32, 39, 33]. In that respect, [46] introduced a batch update algorithm for CaGP and proposed an ELBO objective to facilitate model selection. [15] extended CaGPs by proposing a calibrated probabilistic linear solver, resulting in a calibrated computation-aware GP model.
7 Experimental results
We evaluate the performance of our proposed RCaGP against concurrent baselines on a range of regression datasets, followed by several high-throughput Bayesian Optimization tasks. We conclude by presenting insights gained after carrying out dedicated ablation studies.
GP baselines. We compare RCaGP against RCSVGP [2], SVGP [16], and CaGP [46]. Finally, we set the number of actions for RCaGP and CaGP, and the number of inducing points for SVGP and RCSVGP, to $25$. All baselines are equipped with a Matern $5/2$ kernel and were implemented using GPyTorch [9]. Model hyperparameters are optimized using the ELBO (Equation 13).
Outlier contamination protocol. In our experiments, we follow the settings described by [2]. Unless stated otherwise in dedicated ablation studies, for regression datasets, we sample uniformly at random 10% of the training dataset input-output pairs $({\bf{x}}_{i},y_{i})$, and replace the $y_{i}$’s by asymmetric outliers, i.e., via subtraction of noise sampled from a uniform distribution $\mathcal{U}(3\bar{\sigma},9\bar{\sigma})$, with $\bar{\sigma}$ being the standard deviation of the original observations. For BO tasks, while running the optimization loop, each evaluation has a $25\%$ chance of returning an asymmetric outlier, obtained by addition of noise sampled from a uniform distribution $\mathcal{U}(1\bar{\sigma},2\bar{\sigma})$, with $\bar{\sigma}$ being is the estimated standard deviation of the normalized evaluation function. Finally, for the weight function $w$ (Equation 12), we set the soft threshold $c=Q_{n}(|\mathbf{y}-{\bf{m}}|,1-\epsilon)$ with $\epsilon=0.2$, with $Q_{n}$ the $(1-\epsilon)$-quantile of $|\mathbf{y}-{\bf{m}}|$.
Bayesian Optimization settings. As described during Section 3.3, we integrate the baselines into EULBO, building on the implementation of [28]. We initialize the optimization process for all baselines with $250$ data points sampled uniformly across the search space. All surrogates employ expected improvement (EI) as the acquisition function. Standard BO is applied across all tasks, while trust-region BO (TuRBO [8]) is additionally employed for high-dimensional tasks.
7.1 Regression on UCI datasets
We evaluate GP regression on four UCI datasets with asymmetric outliers (details in Section H.1); results are shown in the left part of Table 1. RCaGP consistently delivers the best performance, combining strong predictive accuracy and well-calibrated uncertainty with acceptable runtimes. Notably, it significantly outperforms both CaGP and RCSVGP in this setting, indicating that the integration of robustness and approximation-awareness is effective.
To test robustness across contamination types, we further consider two additional outlier scenarios, uniform and focused outliers, inspired by [2] (definitions in Section H.3). The corresponding results are shown in Table S1 and Table S2. RCaGP remains the top performer overall, particularly in uniform outliers for MAE, and focused outliers for NLL. These findings support its robustness across contamination types. More importantly, because outlier presence is not always known in advance, we also evaluate performance under clean conditions (right side of Table 1). RCaGP remains competitive and often superior—even without injected noise. One exception is the Energy dataset, where CaGP marginally outperforms RCaGP. This suggests that even in the absence of outliers, the weight function $w$ used by RCaGP (Equation 12) proves useful. An alternative explanation would be that these datasets already contain outliers from the start.
Overall, these results demonstrate that RCaGP is more than the sum of its parts: the combination of RCGP and CaGP is not merely additive, but synergistic, consistently outperforming either component alone—both in outlier-rich and clean regimes.
Table 1: UCI Regression datasets results with asymmetric outliers and without outliers. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits. Bolded results refer to the best baseline. Lower is better.
Asymmetric outliers No outliers
Boston Energy Yacht Parkinsons Boston Energy Yacht Parkinsons
SVGP MAE 0.749 ± 0.062 0.607 ± 0.066 0.837 ± 0.103 0.766 ± 0.090 0.506 ± 0.048 0.407 ± 0.030 0.672 ± 0.059 0.701 ± 0.078
NLL 1.442 ± 0.037 1.298 ± 0.036 1.460 ± 0.086 1.475 ± 0.062 1.335 ± 0.036 1.207 ± 0.017 1.358 ± 0.042 1.407 ± 0.063
clock-time 1.480 ± 0.260 2.770 ± 0.070 0.960 ± 0.070 0.730 ± 0.170 1.570 ± 0.370 2.930 ± 0.310 0.950 ± 0.030 0.520 ± 0.030
CaGP MAE 0.738 ± 0.059 0.562 ± 0.055 0.775 ± 0.078 0.798 ± 0.084 0.488 ± 0.042 0.334 ± 0.024 0.512 ± 0.041 0.674 ± 0.077
NLL 1.3863 ± 0.0357 1.253 ± 0.027 1.395 ± 0.064 1.440 ± 0.068 1.278 ± 0.040 1.106 ± 0.011 1.263 ± 0.042 1.362 ± 0.067
clock-time 2.410 ± 0.050 3.680 ± 0.080 0.720 ± 0.040 1.220 ± 0.080 2.470 ± 0.040 3.710 ± 0.080 0.730 ± 0.060 0.840 ± 0.010
RCSVGP MAE 0.532 ± 0.050 0.566 ± 0.060 0.477 ± 0.049 0.627 ± 0.089 0.574 ± 0.056 0.612 ± 0.034 0.646 ± 0.076 0.648 ± 0.071
NLL 1.298 ± 0.038 1.280 ± 0.034 1.241 ± 0.035 1.346 ± 0.070 1.332 ± 0.048 1.324 ± 0.033 1.322 ± 0.047 1.403 ± 0.076
clock-time 1.500 ± 0.070 2.860 ± 0.060 1.010 ± 0.030 0.750 ± 0.020 1.560 ± 0.040 2.970 ± 0.120 1.010 ± 0.030 0.570 ± 0.010
RCaGP MAE 0.477 ± 0.046 0.380 ± 0.038 0.435 ± 0.052 0.586 ± 0.084 0.428 ± 0.053 0.393 ± 0.024 0.436 ± 0.046 0.538 ± 0.078
NLL 1.272 ± 0.036 1.162 ± 0.012 1.253 ± 0.043 1.317 ± 0.067 1.261 ± 0.049 1.160 ± 0.008 1.251 ± 0.040 1.311 ± 0.070
clock-time 2.450 ± 0.110 3.790 ± 0.130 0.870 ± 0.060 1.250 ± 0.090 2.580 ± 0.150 3.780 ± 0.190 0.910 ± 0.070 0.920 ± 0.020
7.2 High-Throughput Bayesian Optimization
We consider 4 tasks: Hartmann6D, Lunar12D, Rover60D and Lasso-DNA180D (Section H.2). Figure 3 reports the best value found during BO, averaged over 20 trials, with $\pm$1 standard deviation. To ensure fair comparisons despite outlier contamination, we report the uncontaminated mean best value. In the top row (columns 1–4), using EULBO-EI and under asymmetric outliers, RCaGP clearly outperforms all baselines on Lasso DNA 180D, and ranks first or on par with CaGP on Hartmann 6D, Lunar12D, and Rover60D. SVGP and RCSVGP display the same performance. The bottom row (columns 1-4) shows results with Determinantal Point Process (DPP)-based inducing point selection [30], still in the case of asymmetric outliers. RCaGP remains the top performer on DNA and performs similarly to CaGP on other tasks, with RCSVGP slightly ahead on Hartmann. Substituting EI with Thompson Sampling in DPP-BO leads to equal performance across baselines (Figure S1).
Next, Figure S2 presents results with TurBO, a BO variant tailored to high dimensions [8]. RCaGP exhibits strong average performance, and notably, it is the only method that never ranks last in a statistically significant manner. Lastly, as one might not know beforehand whether outliers will affect the trial, column 5 in Figure 3 shows results for EULBO EI on Hartmann and Lunar without outliers. RCaGP maintains a clear advantage on Hartmann, while all baselines perform similarly on Lunar.
Together, these results highlight RCaGP’s versatility and robustness across inducing point allocation strategies (EULBO, DPP), acquisition functions (EI, TS), BO behavior (global BO, trust-region BO), and noise regimes (with or without outliers). Its consistent superiority supports the central claim: combining robustness from RCGP with the uncertainty calibration of CaGP leads to a synergistic improvement—not just an additive one—across a wide range of BO challenges.
Figure 3: High-throughput Bayesian Optimization task. Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std. Columns 1-4 report results under asymmetric outliers contamination, with the $1^{\text{st}}$ row using the EULBO-EI acquisition function, the $2^{\text{nd}}$ row using DPPBO-EI. Column 5 features results without outliers using EULBO-EI.
Figure 4: Ablation studies. (Left) Varying $c$ in weight function $w$ (Equation 12) on UCI regression datasets. MAE and NLL have been normalized for each dataset; the metrics displayed represent the average. (Right) BO results for RCaGP and RCSVGP using expert-driven prior mean function $m$ in $w$. Results are averaged over across 20 repetitions, with $\pm$ 1 standard deviation being shown.
7.3 Ablation studies
Effect of the soft threshold $c$ in the weight function. We evaluate RCaGP’s sensitivity to the threshold $c=Q_{n}(|\mathbf{y}-{\bf{m}}|,1-\epsilon)$, defined as the $(1-\epsilon)$-quantile of $\lvert\mathbf{y}-{\bf{m}}\rvert$ and controlling how many values are treated as outliers, by varying $\epsilon$. Figure 4 (left) shows MAE and NLL averaged over 4 UCI datasets (normalized; per-dataset results in Figure S3). RCaGP consistently achieves lower error across all $\epsilon$, with best results at $0.15$, suggesting treating roughly 15% of the data as potential outliers is optimal. This aligns well with the naturally present and 10% injected asymmetric outliers.
Effect of the mean $m$ in the weight function. Given the central role of the mean function $m$ in $w$ (Equation 12), we compare a constant prior to the expert-driven one from Section 5, assuming perfect outlier correction. This yields four variants: RCaGP/RCSVGP with either constant (const.) or expert (exp.) mean. As shown in Figure 4 (right), the informed prior greatly boosts RCSVGP across BO tasks, and modestly accelerates RCaGP on Lunar12D. See Appendix G for further analysis.
8 Discussion
This paper introduced Robust Computation-Aware Gaussian Processes (RCaGP), a principled framework unifying robust inference and approximation-aware modeling coherently rather than tackling these challenges independently. Our theoretical and empirical results show that RCaGP yields conservative uncertainty estimates, is provably robust to outliers, and consistently outperforms methods addressing only one challenge. Additionally, we proposed an expert-guided prior mean to further enhance robustness in practice. RCaGP represents a meaningful step forward in probabilistic numerics, enabling robust and trustworthy inference for complex, large-scale regression and optimization tasks.
Limitations.
While RCaGP demonstrates strong performance across diverse regression and BO tasks, it inherits certain practical limitations. Like CaGP, it requires full batching, leading to increased GPU memory demands compared to fully stochastic approaches like SVGP. Its performance also depends on manually selecting the projection dimensionality, which cannot be tuned via standard GP model selection techniques. Lastly, broader evaluations across more complex domains remain an avenue for future work.
Supplementary Material - Robust and Computation-Aware Gaussian Processes
The appendix is organized as follows:
  • Appendix A gives the results of additional experiments mentioned in the main text. These include experiments on UCI regression datasets with uniform and focused outliers (Table S1 and Table S2), BO results for DPP-BO-TS (Figure S1), TuRBO (Figure S2), and dataset-specific results for the ablation study conducted in Section 7.3, where $c$ is varied.
  • Appendix B gives the computational complexity of RCaGP.
  • Appendix C formally derives the ELBO for RCaGP introduced in Section 3.2.
  • Appendix D contains the proof for the robustness property presented in Proposition 4.1.
  • Appendix E proves Proposition 4.2, showing that RCaGP’s uncertainty estimates capture the worse-case error over all latent functions.
  • Appendix F contains additional theoretical results: a proposition showing the convergence in mean of RCaGP in RKHS norm (Section F.2), and a proposition showing the lack of robustness of CaGP (Section F.1).
  • Appendix G provides further details regarding the expert-driven prior mean introduced in Section 5 and employed during the ablation study in Section 7.3.
  • Appendix H provides additional details about the UCI regression datasets (Section H.1), the test functions employed in BO (Section H.2), and the outlier-contamination protocols studied in this work (Section H.3).
  • Appendix I provides tables that report the hyperparmeters used for UCI regression and high-throughput BO experiments (Section I.1) and the description of the computing resources used for our experiments (Section I.2).
Appendix A Additional experiments
The results of the UCI regression experiments in the presence of uniform and focused outliers are presented in Table S1 and Table S2, respectively. The MAE and NLL metrics indicate that RCaGP consistently outperforms the baseline methods across most datasets. These findings demonstrate the robustness and versatility of RCaGP in handling different types of outliers. However, this performance improvement comes at the cost of increased computational time compared to the baselines.
The results of Bayesian optimization (BO) experiments using the Thompson Sampling (TS) acquisition function (AF) and Trust-Region BO (TuRBO) are presented in Figure S1 and Figure S2, respectively. In experiments utilizing the TF acquisition function, no model demonstrates a statistically significant performance advantage. Furthermore, the best values identified by these models are generally lower than those obtained with the BO-EI method. We hypothesize that this performance degradation arises from the reliance of the TS approach on posterior sampling, which may reduce the effectiveness of the robustness properties inherent in RCaGP. Subsequent BO experiments using the TuRBO framework show improved performance across all models and tasks—except for the Lunar task. In this case, TuRBO yields lower results for all models. This suggests that the presence of outliers in the Lunar task may impede TuRBO’s ability to effectively explore and identify the global optimum.
The MAE and NLL plots for varying values of $\epsilon$ in the weight function $w$ are shown in Figure S3. Across all datasets, RCaGP consistently yields lower MAE and NLL values than RCSVGP, indicating superior predictive performance. Furthermore, both metrics improve when $\epsilon$ is set to 0.1 or 0.15, but begin to deteriorate once $\epsilon$ exceeds 0.2. This trend suggests that aligning $\epsilon$ with the true proportion of outliers in the data enhances model robustness and overall performance.
Table S1: UCI Regression datasets results in the presence of uniform outliers with $\epsilon=0.1$. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
Boston Energy Yacht Parkinsons
SVGP MAE 0.799 ± 0.054 0.918 ± 0.070 0.899 ± 0.099 0.837 ± 0.093
NLL 1.511 ± 0.047 1.486 ± 0.042 1.558 ± 0.095 1.523 ± 0.057
clock-time 1.580 ± 0.050 2.680 ± 0.100 1.000 ± 0.080 0.800 ± 0.010
CaGP MAE 0.782 ± 0.042 0.908 ± 0.046 0.861 ± 0.094 0.826 ± 0.087
NLL 1.483 ± 0.047 1.477 ± 0.029 1.542 ± 0.098 1.479 ± 0.073
clock-time 2.570 ± 0.040 3.640 ± 0.220 0.810 ± 0.280 0.960 ± 0.040
RCSVGP (const.) MAE 0.763 ± 0.065 0.921 ± 0.082 0.799 ± 0.095 0.8567 ± 0.1017
NLL 1.482 ± 0.066 1.484 ± 0.059 1.496 ± 0.089 1.5494 ± 0.0753
clock-time 1.620 ± 0.040 2.700 ± 0.060 1.050 ± 0.030 0.830 ± 0.020
RCaGP (const.) MAE 0.743 ± 0.053 0.910 ± 0.046 0.762 ± 0.097 0.811 ± 0.077
NLL 1.475 ± 0.062 1.490 ± 0.036 1.533 ± 0.099 1.469 ± 0.075
clock-time 2.630 ± 0.040 3.770 ± 0.110 0.960 ± 0.050 1.080 ± 0.020
Table S2: UCI Regression datasets results in the presence of focused outliers with $\epsilon=0.1$. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
Boston Energy Yacht Parkinsons
SVGP MAE 0.627 ± 0.051 0.337 ± 0.021 0.630 ± 0.053 0.672 ± 0.078
NLL 1.405 ± 0.031 1.239 ± 0.009 1.255 ± 0.057 1.433 ± 0.053
clock-time 1.500 ± 0.050 2.800 ± 0.080 1.000 ± 0.060 0.770 ± 0.020
CaGP MAE 0.677 ± 0.061 0.436 ± 0.051 0.525 ± 0.050 0.750 ± 0.073
NLL 1.425 ± 0.055 1.181 ± 0.03 1.277 ± 0.048 1.462 ± 0.064
clock-time 2.460 ± 0.060 3.670 ± 0.070 0.830 ± 0.290 0.930 ± 0.010
RCSVGP (const.) MAE 0.709 ± 0.100 0.634 ± 0.070 0.666 ± 0.072 0.551 ± 0.079
NLL 1.419 ± 0.088 1.436 ± 0.096 1.321 ± 0.087 1.392 ± 0.068
clock-time 1.550 ± 0.040 2.850 ± 0.070 1.040 ± 0.040 0.790 ± 0.030
RCaGP (const.) MAE 0.505 ± 0.069 0.390 ± 0.023 0.437 ± 0.053 0.600 ± 0.082
NLL 1.323 ± 0.073 1.163 ± 0.009 1.268 ± 0.051 1.374 ± 0.082
clock-time 2.570 ± 0.050 3.780 ± 0.110 0.930 ± 0.060 1.010 ± 0.040
Figure S1: High-throughput Bayesian Optimization task under asymmetric outliers using DPP-BO and the Thompson Sampling (TS) acquisition strategy. Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std.
Figure S2: High-throughput Bayesian Optimization task using TurBO under asymmetric outliers. Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std. Top row: EULBO-TuRBO-EI. Bottom row: DPP-TuRBO-EI.
Figure S3: Dataset-specific results for the ablation study conducted in Section 7.3, varying $c$ in weight function $w$ (Equation 12).
Appendix B Computational Complexity
Assuming a constant mean prior, RCaGP exhibits the same time and memory complexity as CaGP, as both rely on the same batch update algorithm. Note that the weight function requires ${\mathcal{O}}(n)$ for time and space complexity, thus it does not affect the complexity of the algorithm. The time complexity of RCaGP is ${\mathcal{O}}(n\,i\,\max(i,k))$, where $k=n/i$ is the number of non-zero entries for each column ${\bf{S}}_{i}$. In addition, RCaGP and CaGP have linear memory complexity ${\mathcal{O}}(ni)$.
Appendix C Evidence Lower Bound for RCaGP
Proposition C.1.
Define the variational family
${\mathcal{Q}}\triangleq\{q_{i}({\bf{f}})={\mathcal{N}}({\bf{f}};\hat{\mu}_{i}({\bf{X}}),\hat{k}_{i}({\bf{X}},{\bf{X}}))|{\bf{S}}_{i}\in^{n\times i}\}$ (S1)
and robust loss function
$L^{w}_{n}({\bf{f}},{\bf{y}},{\bf{X}})=\frac{1}{2n}\left({\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}-2{\bf{f}}^{\top}\nu+C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})\right),$ (S2)
where $\nu=\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and $C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})={\bf{y}}^{\top}\sigma_{\mathrm{noise}}^{-2}\mathrm{diag}({\bf{w}}^{2}){\bf{y}}-2\nabla_{y}{\bf{y}}^{\top}{\bf{w}}^{2}$. Then, the evidence lower bound of RCaGP is given by
$\ell_{\mathrm{ELBO}}^{\mathrm{RCaGP}}={\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]-{\mathrm{KL}}[q_{i}({\bf{f}})\|p({\bf{f}})]$ (S3)
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}}))+\hat{\mu}_{i}({\bf{X}})^{\top}\nu$
$\quad-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})-\frac{1}{2}(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}+\log\det({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})-i\log({\sigma_{\mathrm{noise}}^{2}})$
$\quad-\log\det({\bf{S}}_{i}^{\top}{\bf{S}}_{i})-\log\det({\bf{J}}_{\bf{w}})-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}))$ (S4)
where $\bar{{\bf{v}}}_{i}=({\bf{S}}^{\top}\tilde{{\bf{K}}}{\bf{S}})^{-1}{\bf{S}}^{\top}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and $p^{w}({\bf{y}}|{\bf{f}})=\exp(-nL^{w}_{n}({\bf{f}},{\bf{y}},{\bf{X}}))$ are the projected representer weights and the pseudo-likelihood, respectively.
Proof:
The ELBO is given by
$\ell_{\mathrm{ELBO}}^{\mathrm{RCaGP}}={\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]-{\mathrm{KL}}[q_{i}({\bf{f}})\|p({\bf{f}})]$ (S5)
We first compute the expected loss function term:
[…]
$=\int-\frac{1}{2}({\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}-2{\bf{f}}^{\top}\nu+C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}}))q_{i}({\bf{f}})d{\bf{f}}$ (S7)
[…]
Next, we apply the identity ${\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}={\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})$ (see [2]):
${\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]$ $=-\frac{1}{2}{\mathbb{E}}_{q_{i}({\bf{f}})}[{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})]+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}^{\top}\nu]-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S10)
Since the expectation of the trace is equal to the trace of the expectation, we find that
${\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]$ $=-\frac{1}{2}{\mathrm{tr}}({\mathbb{E}}_{q_{i}({\bf{f}})}[\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}])+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}^{\top}\nu]-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S11)
$=-\frac{1}{2}{\mathrm{tr}}({\mathbb{E}}_{q_{i}({\bf{f}})}[\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}])+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}]^{\top}\nu\ -\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S12)
By definition of mean and variance, we have that ${\mathbb{E}}_{p({\bf{x}})}[{\bf{x}}{\bf{x}}^{\top}]={\mathbb{V}}_{p({\bf{x}})}[{\bf{x}}]+{\mathbb{E}}_{p({\bf{x}})}[{\bf{x}}]\,{\mathbb{E}}_{p({\bf{x}})}[{\bf{x}}]^{\top}$. Applying this identity, we obtain
${\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]$ $=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}{\bf{f}}^{\top}]\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}]^{\top}\nu\ -\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S13)
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}({\mathbb{V}}_{q_{i}({\bf{f}})}[{\bf{f}}]+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}]{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}]^{\top})\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})$
$\,\quad+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}]^{\top}\nu-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S14)
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}(\hat{k}_{i}({\bf{X}},{\bf{X}})+\hat{\mu}_{i}({\bf{X}})\hat{\mu}_{i}({\bf{X}})^{\top})\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})+\hat{\mu}_{i}({\bf{X}})^{\top}\nu$
$\,\quad-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S15)
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}})\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1/2})$
$\,\quad+\hat{\mu}_{i}({\bf{X}})^{\top}\nu-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S16)
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1}\hat{\mu}_{i}({\bf{X}}))$
$\,\quad+\hat{\mu}_{i}({\bf{X}})^{\top}\nu-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S17)
Next, we compute the KL term, where both $q({\bf{f}})$ and $p({\bf{f}})$ are multivariate Gaussian:
${\mathrm{KL}}[q_{i}({\bf{f}})\,\|\,p({\bf{f}})]$ $=\frac{1}{2}((\hat{\mu}_{i}({\bf{X}})-m({\bf{X}}))^{\top}{\bf{K}}^{-1}(\hat{\mu}_{i}({\bf{X}})-m({\bf{X}}))+\log\left(\frac{\det({\bf{K}})}{\det(\hat{k}_{i}({\bf{X}},{\bf{X}}))}\right)$
$\,\quad+\,{\mathrm{tr}}({\bf{K}}^{-1}\hat{k}_{i}({\bf{X}},{\bf{X}}))-n)$ (S18)
$=\frac{1}{2}(({\bf{K}}\tilde{{\bf{C}}}_{i}({\bf{y}}-m_{w}({\bf{X}})))^{\top}{\bf{K}}^{-1}({\bf{K}}\tilde{{\bf{C}}}_{i}({\bf{y}}-m_{w}({\bf{X}})))-\log\det({\bf{K}}^{-1}\,\hat{k}_{i}({\bf{X}},{\bf{X}}))$
$\,\quad+\,{\mathrm{tr}}({\bf{I}}_{n\times n}-\tilde{{\bf{C}}}_{i}{\bf{K}})-n)$ (S19)
$=\frac{1}{2}(({\bf{y}}-m_{w}({\bf{X}}))^{\top}\tilde{{\bf{C}}}_{i}{\bf{K}}\tilde{{\bf{C}}}_{i}({\bf{y}}-m_{w}({\bf{X}}))-\log\det({\bf{I}}_{n\times n}-\tilde{{\bf{C}}}_{i}{\bf{K}})$
$\quad+{\mathrm{tr}}({\bf{I}}_{n\times n}-\tilde{{\bf{C}}}_{i}{\bf{K}})-n)$ (S20)
$=\frac{1}{2}(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}-\log\det({\bf{I}}_{n\times n}-\tilde{{\bf{C}}}_{i}{\bf{K}})+{\mathrm{tr}}({\bf{I}}_{n\times n}-\tilde{{\bf{C}}}_{i}{\bf{K}})-n)$ (S21)
$=\frac{1}{2}\left(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}-\log\det({\bf{I}}_{n\times n}-\tilde{{\bf{C}}}_{i}{\bf{K}})-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i})\right)$ (S22)
Here, we apply the Weinstein–Aronszajn identity:
$=\frac{1}{2}\left(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}-\log\det({\bf{I}}_{i\times i}-({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i})-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i})\right)$ (S23)
$=\frac{1}{2}\left(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}-\log\det(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i}-{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}))-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i})\right)$ (S24)
$=\frac{1}{2}\left(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}-\log\det(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\sigma_{\mathrm{noise}}^{2}}{\bf{S}}_{i}^{\top}{\bf{J}}_{\bf{w}}{\bf{S}}_{i})-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i})\right)$ (S25)
$=\frac{1}{2}\left(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}+\log\det({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})-\log\det({\sigma_{\mathrm{noise}}^{2}}{\bf{S}}_{i}^{\top}{\bf{J}}_{\bf{w}}{\bf{S}}_{i})-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i})\right)$ (S26)
$=\frac{1}{2}(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}+\log\det({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})-i\log({\sigma_{\mathrm{noise}}^{2}})-\log\det({\bf{J}}_{\bf{w}})-\log\det({\bf{S}}_{i}^{\top}{\bf{S}}_{i})$
$\quad-{\mathrm{tr}}(({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}))$ (S27)
Appendix D Robustness Property
The following lemma contributes to the proof of Proposition 4.1.
Lemma D.1.
For an arbitrary matrice $\hat{{\bf{S}}}\in^{m\times n}$ and positive semidefinite matrice $\hat{{\bf{B}}}\in^{n\times n}$, we have that
$(\hat{{\bf{S}}}\hat{{\bf{B}}}\hat{{\bf{S}}}^{\top})^{-1}$ $=\hat{{\bf{S}}}^{+\top}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{G}}}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{+}$ (S28)
where we define $\hat{{\bf{G}}}={\bf{I}}-\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}})(\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}}))^{+}$ and + denotes the Moore-Penrose inverse.
Proof:
The whole proof is derived from an answer to a question posted on the Mathematics Stack Exchange Forums, which we write here for conciseness.
Denote $\hat{{\bf{O}}}={\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}}$ and ${\bf{H}}(\alpha)=(\hat{{\bf{S}}}(\alpha{\bf{I}}+\hat{{\bf{B}}}^{-1})^{-1}\hat{{\bf{S}}}^{\top})^{-1}$. We also note that
$(\hat{{\bf{S}}}\hat{{\bf{B}}}\hat{{\bf{S}}}^{\top})^{-1}=\lim_{\alpha\rightarrow 0}{\bf{H}}(\alpha)$ (S29)
By applying Woodbury matrix identity, we can rewrite ${\bf{H}}(\alpha)$ as follows:
${\bf{H}}(\alpha)=\left(\frac{1}{\alpha}\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top}-\frac{1}{\alpha}\hat{{\bf{S}}}\hat{{\bf{B}}}^{-1/2}\left({\bf{I}}+\frac{1}{\alpha}\hat{{\bf{B}}}^{-1}\right)^{-1}\frac{1}{\alpha}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{\top}\right)^{-1}$ (S30)
Using the fact that $\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top}$ is invertible and applying the Woodbury matrix identity for the second time, we obtain
${\bf{H}}(\alpha)=\alpha(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}-(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}\hat{{\bf{S}}}\hat{{\bf{B}}}^{-1/2}$
$\quad(-({\bf{I}}+\frac{1}{\alpha}\hat{{\bf{B}}}^{-1})+\frac{1}{\alpha}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{\top}(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}\hat{{\bf{S}}}\hat{{\bf{B}}}^{-1/2})^{-1}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{\top}(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}$ (S31)
$=\alpha(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}+(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}\hat{{\bf{S}}}\hat{{\bf{B}}}^{-1/2}({\bf{I}}+\frac{1}{\alpha}\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{\top}(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}\hat{{\bf{S}}})\hat{{\bf{B}}}^{-1/2})^{-1}$
$\quad\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{\top}(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}$ (S32)
We note that
$\hat{{\bf{S}}}^{\top}(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}=\hat{{\bf{S}}}^{+}$ (S33)
${\bf{I}}-\hat{{\bf{S}}}^{\top}(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}\hat{{\bf{S}}}=\hat{{\bf{O}}}$ (S34)
Then, we rewrite ${\bf{H}}(\alpha)$ as follows:
${\bf{H}}(\alpha)=\alpha(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}+\hat{{\bf{S}}}^{+\top}\hat{{\bf{B}}}^{-1/2}\left({\bf{I}}+\frac{1}{\alpha}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{O}}}\hat{{\bf{O}}}\hat{{\bf{B}}}^{-1/2}\right)^{-1}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{+}$ (S35)
Applying the Woodbury matrix identity for the third time gives us
${\bf{H}}(\alpha)=\alpha(\hat{{\bf{S}}}\hat{{\bf{S}}}^{\top})^{-1}+\hat{{\bf{S}}}^{+\top}\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{B}}}^{-1/2}\hat{{\bf{O}}}(\alpha{\bf{I}}+\hat{{\bf{O}}}\hat{{\bf{B}}}^{-1}\hat{{\bf{O}}})^{-1}\hat{{\bf{O}}}\hat{{\bf{B}}}^{-1/2})\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{+}$ (S36)
Since the Moore-Penrose inverse of a matrice $\mathbf{A}$ is a limit:
$\mathbf{A}^{+}=\lim_{\alpha\rightarrow 0}(\mathbf{A}^{\top}\mathbf{A}+\alpha{\bf{I}})^{-1}\mathbf{A}^{\top}=\lim_{\alpha\rightarrow 0}\mathbf{A}^{\top}(\mathbf{A}\mathbf{A}^{\top}+\alpha{\bf{I}})^{-1}$ (S37)
We can take the limit of ${\bf{H}}(\alpha)$ as $\alpha\rightarrow 0$ and apply the limit relation above to obtain the following result:
$(\hat{{\bf{S}}}\hat{{\bf{B}}}\hat{{\bf{S}}}^{\top})^{-1}=\hat{{\bf{S}}}^{+\top}\hat{{\bf{B}}}^{-1/2}\underbrace{({\bf{I}}-\hat{{\bf{B}}}^{-1/2}\hat{{\bf{O}}}(\hat{{\bf{B}}}^{-1/2}\hat{{\bf{O}}})^{+})}_{\hat{{\bf{G}}}}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{+}$ (S38)
PIF of RCaGP We now prove Proposition 4.1: Let ${\bf{f}}\sim{\mathcal{GP}}(m,k)$ denote the RCaGP prior, and let $i\in{0,\dots,n}$ represent the number of actions in RCaGP. Assume the observation noise is $\varepsilon\sim{\mathcal{N}}(\mathbf{0},{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})$. Define constants $C_{k}^{\prime}\in;k=1,2,3$, which are independent of $y_{m}^{c}$. For any $i$ and assuming $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$, the PIF of RCaGP is given by:
$\mathrm{PIF}_{\mathrm{RCaGP}}(y_{m}^{c},{\mathcal{D}})=C_{2}^{\prime}(w({\bf{x}}_{m},y_{m}^{c})^{2}y_{m}^{c})^{2}+C_{3}^{\prime}.$ (S39)
Therefore, if $\sup_{{\bf{x}},y}y\,w({\bf{x}},y)^{2}<\infty$, RCaGP regression is robust since $\sup_{y_{m}^{c}}|\mathrm{PIF}_{\mathrm{RCaGP}}(y_{m}^{c},{\mathcal{D}})|<\infty$.
Proof:
Without loss of generality, we aim to prove the bound for $m=n$. We can extend the proof for an arbitrary $m\in\{1,\dots,n\}$. Let $p^{w}({\bf{f}}|{\mathcal{D}})={\mathcal{N}}({\bf{f}};\hat{{\boldsymbol{\mu}}}_{i},\hat{{\bf{K}}}_{i})$ and $p^{w}({\bf{f}}|{\mathcal{D}}_{m}^{c})={\mathcal{N}}({\bf{f}};\hat{{\boldsymbol{\mu}}}_{i}^{c},\hat{{\bf{K}}}_{i}^{c})$ be the uncontaminated and contaminated computation-aware RCGP, respectively. Here,
$\hat{{\boldsymbol{\mu}}}_{i}={\bf{m}}+{\bf{K}}\tilde{{\bf{C}}}_{i}\tilde{{\bf{v}}}_{i}$ (S40)
$\hat{{\bf{K}}}_{i}={\bf{K}}\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}}$ (S41)
$\hat{{\boldsymbol{\mu}}}_{i}^{c}={\bf{m}}+{\bf{K}}\tilde{{\bf{C}}}_{i}^{c}\tilde{{\bf{v}}}_{i}^{c}$ (S42)
$\hat{{\bf{K}}}_{i}^{c}={\bf{K}}\tilde{{\bf{C}}}_{i}^{c}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}}$ (S43)
where ${\bf{w}}^{c}=[w({\bf{x}}_{1},y_{1}),\dots,w({\bf{x}}_{n},y_{n}^{c})]^{\top}$. The PIF has the following form
$\mathrm{PIF}_{\mathrm{RCaGP}}(y_{m}^{c},{\mathcal{D}},i)=\frac{1}{2}\left(\underbrace{\mathrm{Tr}((\hat{{\bf{K}}}_{i}^{c})^{-1}\hat{{\bf{K}}}_{i})-n}_{(1)}+\underbrace{(\hat{{\boldsymbol{\mu}}}_{i}^{c}-\hat{{\boldsymbol{\mu}}}_{i})^{\top}(\hat{{\bf{K}}}_{i}^{c})^{-1}(\hat{{\boldsymbol{\mu}}}_{i}^{c}-\hat{{\boldsymbol{\mu}}}_{i})}_{(2)}+\underbrace{\ln\left(\frac{\mathrm{det}(\hat{{\bf{K}}}_{i}^{c})}{\mathrm{det}(\hat{{\bf{K}}}_{i})}\right)}_{(3)}\right)$ (S44)
We first derive the bound for $(1)$:
$(1)$ $=\mathrm{Tr}((\hat{{\bf{K}}}_{i}^{c})^{-1}\hat{{\bf{K}}}_{i})-n$ (S45)
$=\mathrm{Tr}\left(({\bf{K}}\tilde{{\bf{C}}}_{i}^{c}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})^{-1}{\bf{K}}\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}}\right)-n$ (S46)
$=\mathrm{Tr}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1}(\tilde{{\bf{C}}}_{i}^{c})^{-1}\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})-n$ (S47)
$\leq\mathrm{Tr}(\sigma^{-2}_{\mathrm{noise}}{\bf{J}}_{{\bf{w}}^{c}}^{-1}(\tilde{{\bf{C}}}_{i}^{c})^{-1})\mathrm{Tr}(\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})-n$ (S48)
$\leq\mathrm{Tr}(\sigma^{-2}_{\mathrm{noise}}{\bf{J}}_{{\bf{w}}^{c}}^{-1})\mathrm{Tr}((\tilde{{\bf{C}}}_{i}^{c})^{-1})\mathrm{Tr}(\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})-n$ (S49)
The first and second inequality come from the fact that $\mathrm{Tr}(\mathbf{A}\mathbf{F})\leq\mathrm{Tr}(\mathbf{A})\mathrm{Tr}(\mathbf{F})$ for two positive semidefinite matrices $\mathbf{A}$ and $\mathbf{F}$. Since $\mathrm{Tr}(\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})$ does not contain the contamination term, we can write $\bar{C}_{1}=\mathrm{Tr}(\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})$. Let ${\bf{B}}=({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}^{c}{\bf{S}}_{i})^{-1}$ such that ${\bf{C}}_{i}^{c}={\bf{S}}_{i}^{\top}{\bf{B}}{\bf{S}}_{i}^{\top}$. Observe that matrice ${\bf{B}}$ is positive semidefinite. Thus, we can apply Lemma D.1 to obtain the bound of $\mathrm{Tr}((\tilde{{\bf{C}}}_{i}^{c})^{-1})$:
$\mathrm{Tr}((\tilde{{\bf{C}}}_{i}^{c})^{-1})$ $=\mathrm{Tr}(({\bf{S}}_{i}^{\top}{\bf{B}}{\bf{S}}_{i}^{\top})^{-1})$ (S50)
$=\mathrm{Tr}({\bf{S}}_{i}^{+\top}{\bf{B}}^{-1/2}{\bf{G}}{\bf{B}}^{-1/2}{\bf{S}}_{i}^{+})$ (S51)
$\leq\mathrm{Tr}({\bf{S}}_{i}^{+}{\bf{S}}_{i}^{+\top})\mathrm{Tr}({\bf{B}}^{-1/2}{\bf{B}}^{-1/2})\mathrm{Tr}({\bf{G}})$ (S52)
Furthermore, we derive the bound of $\mathrm{Tr}({\bf{G}})$ as follows:
$\mathrm{Tr}({\bf{G}})$ $=\mathrm{Tr}({\bf{I}}-{\bf{B}}^{-1/2}({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})({\bf{B}}^{-1/2}({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i}))^{+})$ (S53)
$=n-\mathrm{Tr}({\bf{B}}^{-1/2}({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})^{+}{\bf{B}}^{-1/2+})$ (S54)
$\leq n-\mathrm{Tr}({\bf{B}}^{-1/2+}{\bf{B}}^{-1/2})\mathrm{Tr}(({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})^{+})$ (S55)
The inequality S52 stems from the trace circular property and the properties of the product of two positive semidefinite matrices. Notably, we observe that $\mathrm{Tr}({\bf{G}})\leq n$ since ${\bf{B}}^{-1/2+}{\bf{B}}^{-1/2}$ and $({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})({\bf{I}}-{\bf{S}}_{i}^{+}{\bf{S}}_{i})^{+}$ in S55 are positive semidefinite matrices. By definition, the trace of a positive semidefinite matrix is non-negative. We then apply the trace circular property for the second time to obtain
$\mathrm{Tr}((\tilde{{\bf{C}}}_{i}^{c})^{-1})$ $\leq n\mathrm{Tr}({\bf{S}}_{i}^{+}{\bf{S}}_{i}^{+\top})\mathrm{Tr}({\bf{B}}^{-1})$ (S56)
$\leq n\mathrm{Tr}({\bf{S}}_{i}^{+}{\bf{S}}_{i}^{+\top})\mathrm{Tr}({\bf{S}}_{i}{\bf{S}}_{i}^{\top})\mathrm{Tr}(\tilde{{\bf{K}}}^{c})$ (S57)
$=\bar{C}_{2}\mathrm{Tr}({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})$ (S58)
where we define $\bar{C}_{2}=n\mathrm{Tr}({\bf{S}}_{i}^{+}{\bf{S}}_{i}^{+\top})\mathrm{Tr}({\bf{S}}_{i}{\bf{S}}_{i}^{\top})$. We then plug S58 into S49 to obtain
$(1)$ $\leq\mathrm{Tr}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1})\mathrm{Tr}({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})\bar{C}_{1}\bar{C}_{2}-n$ (S59)
$=\left(\sum_{j=1}^{n}\left(4\sigma_{\mathrm{noise}}^{-2}w^{2}({\bf{x}}_{j},y_{j})\right)\sum_{k=1}^{n}\left({\bf{K}}_{kk}+\sigma^{4}_{\mathrm{noise}}/2\;w^{-2}({\bf{x}}_{k},y_{k})\right)\right)\bar{C}_{1}\bar{C}_{2}-n$ (S60)
[…]
Next, we derive the bound for the term (2). Following [2], we have that
$(2)\leq\lambda_{\max}((\hat{{\bf{K}}}_{i}^{c})^{-1})\|\hat{{\boldsymbol{\mu}}}_{i}^{c}-\hat{{\boldsymbol{\mu}}}_{i}\|_{1}^{2},$ (S62)
where $\lambda_{\mathrm{max}}((\hat{{\bf{K}}}_{i}^{c})^{-1})$ is the maximum eigenvalue of $(\hat{{\bf{K}}}_{i}^{c})^{-1}$. We expand $\lambda_{\max}((\hat{{\bf{K}}}_{i}^{c})^{-1})$ to derive the following bound:
$\lambda_{\max}((\hat{{\bf{K}}}_{i}^{c})^{-1})$ $=\lambda_{\max}(\sigma^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1}(\tilde{{\bf{C}}}_{i}^{c})^{-1}{\bf{K}}^{-1})$ (S63)
$\leq\lambda_{\max}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1})\lambda_{\max}((\tilde{{\bf{C}}}_{i}^{c})^{-1})\lambda_{\max}({\bf{K}}^{-1})$ (S64)
$=\lambda_{\max}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1})\lambda_{\min}(\tilde{{\bf{C}}}_{i}^{c})\lambda_{\max}({\bf{K}}^{-1})$ (S65)
$\leq\lambda_{\max}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1})\lambda_{\min}((\tilde{{\bf{K}}}^{c})^{-1})\lambda_{\max}({\bf{K}}^{-1})$ (S66)
$\leq\lambda_{\max}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1})(\lambda_{\max}({\bf{K}})+\lambda_{\max}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}}))\lambda_{\max}({\bf{K}}^{-1})$ (S67)
The first inequality follows from the property of the maximum eigenvalue of the product of two positive semidefinite matrices. Subsequently, the second equality is derived from the fact that a matrice’s maximum eigenvalue equals its inverse’s minimum eigenvalue. By the definition $\tilde{{\bf{C}}}_{i}^{c}=(\tilde{{\bf{K}}}^{c})^{-1}-\tilde{{\boldsymbol{\Sigma}}}_{i}^{c}$ and the fact that $\tilde{{\bf{C}}}_{i}^{c}$, $(\tilde{{\bf{K}}}^{c})^{-1}$ and ${\boldsymbol{\Sigma}}_{i}$ are positive semidefinite matrices, the minimum eigenvalue of $\tilde{{\bf{K}}}^{c}$ is less or equal than of $\tilde{{\bf{C}}}_{i}^{c}$, resulting in second inequality. The last inequality follows from the equivalence of the maximum eigenvalue and the addition property of the maximum eigenvalue of two positive semidefinite matrices.
Since ${\bf{J}}_{{\bf{w}}^{c}}^{-1}=\mathrm{diag}(\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}({\bf{w}}^{c})^{2})$, and $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$, it holds that $\lambda_{\max}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}^{c}}^{-1})=\bar{C}_{4}<+\infty$ and $\lambda_{\max}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})=\bar{C}_{5}<+\infty$, such that
$\lambda_{\max}((\hat{{\bf{K}}}_{i}^{c})^{-1})\leq\bar{C}_{4}(\lambda_{\max}({\bf{K}})+\bar{C}_{5})\lambda_{\max}({\bf{K}}^{-1})=\bar{C}_{6}$ (S68)
We substitute $\bar{C}_{6}$ into (2) to obtain
$(2)\leq\bar{C}_{6}\|\hat{{\boldsymbol{\mu}}}_{i}^{c}-\hat{{\boldsymbol{\mu}}}_{i}\|_{1}^{2}$ (S69)
$=\bar{C}_{6}\,\|({\bf{m}}+{\bf{K}}\tilde{{\bf{v}}}_{i}^{c})-({\bf{m}}+{\bf{K}}\tilde{{\bf{v}}}_{i})\|_{1}^{2}$ (S70)
$=\bar{C}_{6}\,\|{\bf{K}}(\tilde{{\bf{C}}}_{i}^{c}({\bf{y}}^{c}-{\bf{m}}_{{\bf{w}}^{c}})-\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{{\bf{w}}}))\|_{1}^{2}$ (S71)
$\leq\bar{C}_{6}\,\|{\bf{K}}\|_{F}\,\|\tilde{{\bf{C}}}_{i}^{c}({\bf{y}}^{c}-{\bf{m}}_{{\bf{w}}^{c}})-\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{{\bf{w}}})\|_{1}^{2}$ (S72)
$\leq q\,\bar{C}_{6}\,\|{\bf{K}}\|_{F}\,\|(\tilde{{\bf{K}}}^{c})^{-1}({\bf{y}}^{c}-{\bf{m}}_{{\bf{w}}^{c}})-(\tilde{{\bf{K}}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})\|_{1}^{2}$ (S73)
$=q\,\bar{C}_{6}\,\|{\bf{K}}\|_{F}\,\|({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})^{-1}({\bf{y}}^{c}-{\bf{m}}_{{\bf{w}}^{c}})-({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})\|_{1}^{2}$ (S74)
for a constant $q<+\infty$. The second equality comes from the definition of $\tilde{{\bf{v}}}_{i}$, and the fourth row follows the Cauchy-Schwarz inequality. Finally, the last inequality holds since $\tilde{{\bf{K}}}_{i}=(\tilde{{\bf{C}}}_{i}^{-1}-\tilde{{\boldsymbol{\Sigma}}}_{i})$. Applying results from [2], we obtain
$(2)$ $\leq\,q\,\bar{C}_{6}\,\|{\bf{K}}\|_{F}\,\|({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})^{-1}({\bf{y}}-{\bf{m}}_{{\bf{w}}^{c}})-({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})({\bf{y}}-{\bf{m}}_{\bf{w}})\|_{1}^{2}$ (S75)
[…]
$\leq\,\bar{C}_{11}+\bar{C}_{12}(w(x_{n},y_{n}^{c})^{2}y_{n}^{c})^{2}$ (S77)
where $\bar{C}_{11}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{7}+\bar{C}_{8})^{2}$ and $\bar{C}_{12}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{9}+\bar{C}_{10})^{2}$. The terms $\bar{C}_{7},\bar{C}_{8},\bar{C}_{9},\bar{C}_{10}$ correspond to $\tilde{C}_{6},\tilde{C}_{8},\tilde{C}_{7},\tilde{C}_{9}$ in [2].
The term (3) can be written as follows:
$(3)$ $=\ln\left(\frac{\mathrm{det}(\hat{{\bf{K}}}_{i}^{c})}{\mathrm{det}(\hat{{\bf{K}}}_{i})}\right)$ (S78)
$=\ln\left(\frac{\mathrm{det}(\tilde{{\bf{C}}}_{i}^{c}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}{\mathrm{det}(\tilde{{\bf{C}}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}})}\right)$ (S79)
$=\ln(\mathrm{det}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}}^{-1}\tilde{{\bf{C}}}_{i}^{-1})\mathrm{det}(\tilde{{\bf{C}}}_{i}^{c})\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}}))$ (S80)
Observe that we can write $\bar{C}_{13}=\ln(\mathrm{det}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{{\bf{w}}}^{-1}\tilde{{\bf{C}}}_{i}^{-1}))$ since it does not contain the contimation term. Furthermore, we obtain
$(3)$ $=\ln(\bar{C}_{13}\,\mathrm{det}(\tilde{{\bf{C}}}_{i}^{c})\,\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}}))$ (S81)
$\leq\ln(\bar{C}_{13}\,\mathrm{det}((\tilde{{\bf{K}}}^{c})^{-1})\det({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}}))$ (S82)
$=\ln\left(\bar{C}_{13}\frac{\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}{\mathrm{det}({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}\right)$ (S83)
$\leq\ln\left(\bar{C}_{13}\frac{\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}{\mathrm{det}({\bf{K}})+\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}\right)$ (S84)
The first inequality follows from the determinant property of positive semidefinite matrices. The last inequality leverages the fact that $\mathrm{det}(\mathbf{A}+\mathbf{F})\geq\mathrm{det}(\mathbf{A})+\mathrm{det}(\mathbf{F})$ for $\mathbf{A}$ and $\mathbf{F}$ are positive semidefinite matrices. Since $\mathrm{det}({\bf{K}}),\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})\geq 0$, we find that
$\ln\left(\frac{\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}{\mathrm{det}({\bf{K}})+\mathrm{det}({\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})}\right)\leq 1,$ (S85)
leading to the following inequality:
$(3)\leq\ln(\bar{C}_{13})=\bar{C}_{14}.$ (S86)
Finally, putting the three terms together, we obtain the following bound:
$\mathrm{PIF}_{\mathrm{RCaGP}}(y_{m}^{c},{\mathcal{D}},i)$ $\leq\bar{C}_{3}+\bar{C}_{11}+\bar{C}_{12}(w(x_{n},y_{n}^{c})^{2}y_{n}^{c})^{2}+\bar{C}_{14}$ (S87)
$=C_{1}^{\prime}(w(x_{n},y_{n}^{c})^{2}y_{n}^{c})^{2}+C_{2}^{\prime}$ (S88)
where $C_{1}^{\prime}=\bar{C}_{12}$ and $C_{2}^{\prime}=\bar{C}_{3}+\bar{C}_{11}+\bar{C}_{14}$.
Appendix E RCaGP Capture the Worst-case Errors
See 4.2
Proof:
Let $c_{j}=[\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})]_{j}$ for $j=1,\dots,n$, where we define $k^{w}(.,.)=k(.,.)+\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}\delta_{w}(.,.)$:
$\delta_{w}({\bf{x}},{\bf{x}}^{\prime})=\begin{array}[]{cc}&\begin{cases}w^{-2}({\bf{x}},y)&{\bf{x}}={\bf{x}}^{\prime}\,\text{and}\,{\bf{x}}\in{\mathcal{D}}\\ 2&{\bf{x}}={\bf{x}}^{\prime}\,\text{and}\,{\bf{x}}\notin{\mathcal{D}}\\ 0&{\bf{x}}\neq{\bf{x}}^{\prime}\end{cases}\end{array},$ (S90)
where ${\mathcal{D}}$ denotes the observation set. Then, applying [20, Lemma 3.9] provides
$\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}h({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}})\right)^{2}=\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}h({\bf{x}})-\sum_{j=1}^{n}c_{j}h({\bf{x}}_{j})\right)^{2}$ (S91)
$=\|k^{w}(.,{\bf{x}})-k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},.)\|^{2}_{{\mathcal{H}}_{k^{w}}}$ (S92)
$=\langle k^{w}(.,{\bf{x}}),k^{w}(.,{\bf{x}})\rangle_{{\mathcal{H}}_{k}^{w}}-2\langle k^{w}(.,{\bf{x}}),k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},.)\rangle_{{\mathcal{H}}_{k}^{w}}$
$\quad+\langle k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},.),k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},.)\rangle_{{\mathcal{H}}_{k}^{w}}$ (S93)
By reproducing property, we have
$\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}h({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}})\right)^{2}=k^{w}({\bf{x}},{\bf{x}})-2k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})+k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{X}})\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})$ (S94)
since ${\bf{x}}\neq{\bf{x}}_{j}$, it holds that $k^{w}({\bf{x}},{\bf{X}})=k({\bf{x}},{\bf{X}})$. By definition, we have $k^{w}({\bf{X}},{\bf{X}})=\tilde{{\bf{K}}}={\bf{K}}+\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}{\bf{J}}_{\bf{w}}$, and following in [45, Eq. (S42)], it holds that $\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}=\tilde{{\bf{C}}_{i}}$. Therefore, we obtain
$\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}h({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}})\right)^{2}$ $=k({\bf{x}},{\bf{x}})+{\sigma_{\mathrm{noise}}^{2}}-2k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k({\bf{X}},{\bf{x}})+k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k({\bf{X}},{\bf{x}})$ (S95)
$=k({\bf{x}},{\bf{x}})+{\sigma_{\mathrm{noise}}^{2}}-k({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}k({\bf{X}},{\bf{x}})$ (S96)
$=\hat{k}_{i}({\bf{x}},{\bf{x}})+{\sigma_{\mathrm{noise}}^{2}}$ (S97)
For the last result, we analogously choose $c_{j}=[(\tilde{{\bf{K}}}^{-1}-\tilde{{\bf{C}}}_{i})k^{w}({\bf{X}},{\bf{x}})]_{j}$. Then, we obtain
$\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}\hat{\mu}^{g}({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}})\right)^{2}=\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}\sum_{j=0}^{n}c_{j}h({\bf{x}}_{j})\right)^{2}$ (S98)
$=\|k({\bf{x}},{\bf{X}})(\tilde{{\bf{K}}}^{-1}-\tilde{{\bf{C}}}_{i})k^{w}({\bf{X}},.)\|^{2}_{{\mathcal{H}}_{k^{w}}}$ (S99)
$=k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{K}}}^{-1}\tilde{{\bf{K}}}\tilde{{\bf{K}}}^{-1}k^{w}({\bf{X}},{\bf{x}})-2k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{K}}}^{-1}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})+$
$\quad\;k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})$ (S100)
$=k({\bf{x}},{\bf{X}})(\tilde{{\bf{K}}}^{-1}-\tilde{{\bf{C}}}_{i})k({\bf{X}},{\bf{x}})$ (S101)
$=\tilde{\sigma}^{2}_{i}({\bf{x}},{\bf{x}})$ (S102)
The consequence of Proposition 4.2 is then formalized through the following corollary:
Corollary E.1.
Assume the conditions of Proposition 4.2 hold. Let $\mu_{\ast}$ be the corresponding mathematical RCGP posterior mean and $\hat{\mu}_{i}$ the RCaGP posterior mean. Then, it holds that
$\frac{|h({\bf{x}})-\hat{\mu}_{i}({\bf{x}})|}{\|h\|_{{\mathcal{H}}_{k^{w}}}}\leq\sqrt{\hat{k}_{i}({\bf{x}},{\bf{x}})+{\sigma_{\mathrm{noise}}^{2}}}$ (S103)
$\frac{\hat{\mu}({\bf{x}})-\hat{\mu}_{i}({\bf{x}})}{\|h\|_{{\mathcal{H}}_{k^{w}}}}\leq\sqrt{\tilde{\sigma}^{2}_{i}({\bf{x}},{\bf{x}})}$ (S104)
The proof follows immediately from Proposition 4.2 and notice that $h/\|h\|_{{\mathcal{H}}_{k^{w}}}$ has unit norm. Corollary E.1 implies that the computational uncertainty provides the pointwise bound relative to RCGP’s mathematical posterior, and the combined uncertainty is the pointwise bound relative to the shifted latent function.
Appendix F Additional Theoretical Results
F.1 CaGP Lacks Robustness
PIF for the CaGP. We aim to prove the following statement: CaGP regression has the PIF for some constant $C_{3}^{\prime}\in$.
$\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)=C_{3}^{\prime}(y_{m}-y_{m}^{c})^{2}$ (S105)
and is not robust: $\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)\rightarrow\infty$ as $|y_{m}^{c}|\rightarrow\infty$.
Proof:
Let $p({\bf{f}}|{\mathcal{D}})={\mathcal{N}}({\bf{f}};{\boldsymbol{\mu}}_{i},{\bf{K}}_{i})$ and $p({\bf{f}}|{\mathcal{D}}_{m}^{c})={\mathcal{N}}({\bf{f}};{\boldsymbol{\mu}}_{i}^{c},{\bf{K}}_{i}^{c})$ be the uncontaminated and contaminated computation-aware GP, respectively. Here,
${\boldsymbol{\mu}}_{i}={\bf{m}}+{\bf{K}}{\bf{v}}_{i}$ (S106)
${\bf{K}}_{i}={\bf{K}}{\bf{C}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{I}}_{n}$ (S107)
${\boldsymbol{\mu}}_{i}^{c}={\bf{m}}+{\bf{K}}{\bf{v}}_{i}^{c}$ (S108)
${\bf{K}}_{i}^{c}={\bf{K}}{\bf{C}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{I}}_{n}$ (S109)
Note that both ${\bf{K}}_{i}$ and ${\bf{K}}_{i}^{c}$ share the same matrice ${\bf{C}}_{i}$. Then, the PIF has the following form:
$\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)=\frac{1}{2}\left(\mathrm{Tr}({\bf{K}}_{i}^{c}{\bf{K}}_{i})-n+({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})^{\top}({\bf{K}}_{i}^{c})^{-1}({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})+\ln\left(\frac{\mathrm{det}({\bf{K}}_{i}^{c})}{\mathrm{det}({\bf{K}}_{i})}\right)\right)$ (S110)
Based on [2], the PIF leads to the following form:
$\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)=\frac{1}{2}\left(({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})^{\top}({\bf{K}}_{i}^{c})^{-1}({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})\right)$ (S111)
Notice that the term ${\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i}$ can be written as
${\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i}$ $=({\bf{m}}+{\bf{K}}{\bf{v}}_{i}^{c})-({\bf{m}}+{\bf{K}}{\bf{v}}_{i})$ (S112)
$={\bf{K}}({\bf{v}}_{i}^{c}-{\bf{v}}_{i})$ (S113)
$={\bf{K}}({\bf{C}}_{i}({\bf{y}}^{c}-{\bf{m}})-{\bf{C}}_{i}({\bf{y}}-{\bf{m}}))$ (S114)
$={\bf{K}}({\bf{C}}_{i}({\bf{y}}^{c}-{\bf{y}}))$ (S115)
Substituting the RHS of Eq. (S115) to ${\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i}$ in Eq. (S111), we obtain
$\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)$ $=\frac{1}{2}({\bf{C}}_{i}({\bf{y}}^{c}-{\bf{y}}))^{\top}{\bf{K}}\left({\bf{K}}{\bf{C}}_{i}{\sigma_{\mathrm{noise}}^{2}}{\bf{I}}\right)^{-1}{\bf{K}}({\bf{C}}_{i}({\bf{y}}^{c}-{\bf{y}}))$ (S116)
$=\frac{1}{2}\sigma^{-2}_{\mathrm{noise}}({\bf{y}}^{c}-{\bf{y}})^{\top}{\bf{C}}_{i}^{\top}{\bf{K}}({\bf{y}}^{c}-{\bf{y}})$ (S117)
Note that ${\bf{y}}$ and ${\bf{y}}^{c}$ have only one exception for the $m-$th element. Thus, we have
$\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)=\frac{1}{2}[{\bf{C}}_{i}^{\top}{\bf{K}}\sigma^{-2}{\bf{I}}]_{mm}(y_{m}^{c}-y_{m})^{2}$ (S118)
F.2 Mean convergence of RCaGP
F.2.1 Empirical-risk minimization problem of RCGP.
We first show the corresponding empirical-risk minimization problem of RCGP whenever ${\bf{m}}=\mathbf{0}$. Following [2, proof of Proposition 3.1], we can rewrite $L_{n}^{w}$ and formulate the RCGP objective as follows:
$\hat{{\bf{f}}}=\mathrm{argmin}_{{\bf{f}}\in{\mathcal{H}}_{k}}\frac{1}{2n}\left(\underbrace{{\bf{f}}^{\top}\lambda^{-1}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}-2{\bf{f}}^{\top}\lambda^{-1}{\bf{J}}_{\bf{w}}^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})}_{L^{w}_{n}}\right)+\frac{1}{2}\|{\bf{f}}\|^{2}_{{\mathcal{H}}_{k}}$ (S119)
for $\lambda>0$. The regularization constant $\lambda$ controls the smoothness of the estimator to avoid overfitting, where the larger the $\lambda$ is, the smoother the resulting estimator $\hat{{\bf{f}}}$ becomes. Next, we show the unique solution to S119 through the following lemma:
Lemma F.1.
If $\lambda>0$ and the kernel $k$ is invertible, the solution to S119 is unique, and is given by
$\hat{f}({\bf{x}})={\bf{k}}_{\bf{x}}({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})=\sum_{j=1}^{n}\hat{\alpha}_{j}k({\bf{x}},{\bf{x}}_{j}),{\bf{x}}\in\mathcal{X}$ (S120)
where
$(\hat{\alpha}_{i},\dots,\hat{\alpha}_{n})=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})\in^{n}$ (S121)
Proof:
The optimization problem in S119 allows us to apply the representer theorem [36]. It implies that the solution of S119 can be written as a weighted sum, i.e.,
$\hat{{\bf{f}}}=\sum_{j=1}^{n}\hat{\alpha}_{j}k(.,{\bf{x}}_{j})$ (S122)
for $\hat{\alpha}_{1},\dots,\hat{\alpha}_{n}\in$. Let $\hat{{\boldsymbol{\alpha}}}=[\hat{\alpha}_{1},\dots,\hat{\alpha}_{n}]^{\top}\in^{n}$. Substituting S122 into S119 provides
$\mathrm{argmin}_{\hat{{\boldsymbol{\alpha}}}\in^{n}}$ $\frac{1}{2n}(\lambda^{-1}\hat{{\boldsymbol{\alpha}}}^{\top}{\bf{K}}{\bf{J}}_{\bf{w}}^{-1}{\bf{K}}\hat{{\boldsymbol{\alpha}}}-2\lambda^{-1}\hat{{\boldsymbol{\alpha}}}^{\top}{\bf{K}}{\bf{J}}_{\bf{w}}^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}}))+\frac{1}{2}\hat{{\boldsymbol{\alpha}}}^{\top}{\bf{K}}\hat{{\boldsymbol{\alpha}}}$ (S123)
Taking the differentiation of the objective w.r.t. ${\boldsymbol{\alpha}}$, setting it equal to zero, and arranging the result yields the following equation:
${\bf{K}}({\bf{K}}+n\,\lambda\,{\bf{J}}_{\bf{w}})\hat{{\boldsymbol{\alpha}}}={\bf{K}}({\bf{y}}-{\bf{m}}_{\bf{w}})$ (S124)
Since the objective in S123 is a convex function of ${\boldsymbol{\alpha}}$, we find that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the minimum of the objective (S119 and S123). Furthermore, we can verify that $L_{n}^{w}$ is a convex function w.r.t. ${\bf{f}}$. Therefore, we conclude that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the unique solution to S119. As a remark, Proposition F.1 closely connects with [20, Theorem 3.4].
F.2.2 Proof of Proposition F.3
Relative bound errors. We first provide the equivalence of [45, Proposition 2]:
Proposition F.2.
For any choice of actions a relative bound error $\hat{\rho}(i)$ s.t. $\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|_{\tilde{{\bf{K}}}}\leq\hat{\rho}(i)\|\hat{{\bf{v}}}\|_{\tilde{{\bf{K}}}}$ is given by
$\hat{\rho}(i)=(\bar{{\bf{v}}}^{\top}({\bf{I}}-\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}})\bar{{\bf{v}}})^{1/2}\leq\lambda_{\max}({\bf{I}}-\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}})\leq 1$ (S125)
where $\bar{{\bf{v}}}=\hat{{\bf{v}}}/\|\tilde{{\bf{v}}}\|_{\tilde{{\bf{K}}}}$.
The proof is by directly substituting ${\bf{C}}_{i},\hat{{\bf{K}}},{\bf{v}}_{\ast}$ in [45] with $\tilde{{\bf{C}}}_{i},\tilde{{\bf{K}}},\hat{{\bf{v}}}$, respectively. Next, we establish the convergence theorem of RCaGP through the following proposition:
Proposition F.3.
Let ${\mathcal{H}}_{k}$ be the RKHS w.r.t. kernel $k$, ${\sigma_{\mathrm{noise}}^{2}}>0$ and let $\hat{{\boldsymbol{\mu}}}_{\ast}-{\bf{m}}\in{\mathcal{H}}_{k}$ be the unique solution to following empirical risk minimization problem
$\mathrm{argmin}_{f\in\mathcal{H}_{k}}\frac{1}{2n}\left(\underbrace{{\bf{f}}^{\top}\lambda^{-1}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}-2{\bf{f}}^{\top}\lambda^{-1}{\bf{J}}_{\bf{w}}^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})}_{L^{w}_{n}}\right)+\frac{1}{2}\|{\bf{f}}\|^{2}_{{\mathcal{H}}_{k}}$ (S126)
which is equivalent to the mathematical RCGP mean posterior shifted by prior mean ${\bf{m}}$. Then for the number of actions $i\in\{0,\dots,n\}$ the RCaGP posterior mean $\hat{{\boldsymbol{\mu}}}_{i}$ satisfies:
$\|\hat{{\boldsymbol{\mu}}}_{\ast}-\hat{{\boldsymbol{\mu}}}_{i}\|_{\mathcal{H}_{k}}\leq\hat{\rho}(i)\,c({\bf{J}}_{\bf{w}})\,\|\hat{{\boldsymbol{\mu}}}_{\ast}-{\bf{m}}\|_{\mathcal{H}_{k}}$ (S127)
where $\hat{\rho}$ is the relative bound errors corresponding to the number of actions $i$ and the constant $c({\bf{J}}_{\bf{w}})=\sqrt{1+\frac{\lambda_{\max}({\bf{J}}_{{\bf{w}}})}{\lambda_{\min}({\bf{K}})}}\rightarrow 1$ as $\lambda_{\max}({\bf{J}}_{{\bf{w}}})\rightarrow 0$.
Proof:
Lemma F.1 implies there exists a unique solution to the corresponding RCGP risk minimization problem. Choosing $\hat{\rho}(i)$ as described in Proposition F.2, we have that $\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|^{2}_{\tilde{{\bf{K}}}}\leq\hat{\rho}(i)\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\tilde{{\bf{K}}}}$, where $\tilde{{\bf{v}}}_{0}=\mathbf{0}$. Then, for $i\in\{0,\dots,n\}$ we find that
$\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|_{\bf{K}}^{2}$ $\leq\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|^{2}_{\tilde{{\bf{K}}}}\leq\hat{\rho}^{2}(i)\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\tilde{{\bf{K}}}}^{2}$ (S128)
$\leq\hat{\rho}(i)^{2}\left(\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\bf{K}}^{2}+\frac{\lambda_{\max}({\bf{J}}_{\bf{w}})}{\lambda_{\min}({\bf{K}})}\lambda_{\min}({\bf{K}})\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{2}^{2}\right)$ (S129)
$\leq\hat{\rho}(i)^{2}\left(\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\bf{K}}^{2}+\frac{\lambda_{\max}({\bf{J}}_{\bf{w}})}{\lambda_{\min}({\bf{K}})}\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\bf{K}}^{2}\right)$ (S130)
$\leq\hat{\rho}(i)^{2}\left(1+\frac{\lambda_{\max}({\bf{J}}_{{\bf{w}}})}{\lambda_{\min}({\bf{K}})}\right)\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\bf{K}}^{2}$ (S131)
The third inequality stems from the definition of ${\bf{J}}_{\bf{w}}$ and the fact that the maximum eigenvalue of a diagonal matrix is the largest component of its diagonal. Assuming $\hat{\mu}_{i}(.)=m(.)+\sum_{j=1}^{n}(\tilde{{\bf{v}}}_{i})_{j}k(.,{\bf{x}}_{j})=m(.)+k(.,{\bf{X}})\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and applying result from [45], we have that
$\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|_{\bf{K}}^{2}=\|\hat{{\boldsymbol{\mu}}}_{\ast}-\hat{{\boldsymbol{\mu}}}_{i}\|^{2}_{{\mathcal{H}}_{k}}$ (S132)
Combining both results and defining $c({\bf{J}}_{\bf{w}})=\left(1+\frac{\lambda_{\max}({\bf{J}}_{{\bf{w}}})}{\lambda_{\min}({\bf{K}})}\right)$, we obtain
$\|\hat{{\boldsymbol{\mu}}}_{\ast}-\hat{{\boldsymbol{\mu}}}_{i}\|_{{\mathcal{H}}_{k}}=\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|_{\bf{K}}\leq\hat{\rho}(i)c({\bf{J}}_{\bf{w}})\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\bf{K}}=\hat{\rho}(i)c({\bf{J}}_{\bf{w}})\|\hat{{\boldsymbol{\mu}}}_{\ast}-{\bf{m}}\|_{{\mathcal{H}}_{k}}$ (S133)
Appendix G Details of Expert-guided Robust Mean Prior
In this section, we provide the posterior inference details of our probabilistic user model for defining the informative mean prior. We first discuss the posterior inference of latent variables $\bar{\delta}_{o}$, governing human expert decisions in identifying outliers. For every outlier candidate $({\bf{x}}_{o},\hat{y}_{o})$, we have that
[…]
$p(\bar{\delta}_{o}|\bar{o}_{o})\propto p(\bar{\delta}_{o})\,p(\bar{o}_{o}|\bar{\delta}_{o})$ posterior
$\qquad\quad\;\;=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})\left(\bar{\delta}_{o}(\bar{o}_{o})+(1-\bar{\delta}_{o})(1-\bar{o}_{o})\right)$
$\qquad\quad\;\;=\mathcal{B}(\alpha=\alpha_{o}+\bar{o}_{o},\beta=\beta_{o}+1-\bar{o}_{o}),$ (S136)
where we define the hyperparameters $\alpha_{0}$ and $\beta_{0}$ as follows:
$\alpha_{o}=|z_{o}|=\left|\frac{\hat{y}_{o}-\bar{\mu}}{\bar{\sigma}}\right|\qquad\text{and}\qquad\beta_{o}\approx 0.$
The posterior $p(\bar{\delta}_{o}|\bar{o}_{o})$ exhibits conjugacy, allowing us to derive a closed-form expression, i.e., Beta distribution. In this framework, we define $\alpha_{o}$ using the z-score of the outlier candidates. The z-score information effectively recognizes outliers by comparing each data point to the sample mean and standard deviation. Next, we set $\beta_{o}$ close to zero, reflecting that the expert is confident that $\bar{y}_{o}$ is an outlier. Given the posterior $p(\bar{\delta}_{o}|\bar{o}_{o})$, the expected value ${\mathbb{E}}_{p(\bar{\delta}_{o}|\bar{o}_{o})}[\bar{\delta}_{o}]$ is expressed by:
${\mathbb{E}}_{p(\bar{\delta}_{o}|\bar{o}_{o})}[\bar{\delta}_{o}]=\frac{\alpha}{\alpha+\beta}=\frac{\alpha_{o}+\bar{o}_{o}}{\alpha_{o}+\beta_{o}+1}$ (S137)
Next, we outline the posterior inference details for the outlier corrections provided by the human expert:
[…]
$p(\bar{\mu}_{o}|\bar{y}_{o})\propto p(\bar{\mu}_{o})\,p(\bar{y}_{o}|\bar{\mu}_{o})$ posterior (S140)
$\qquad\quad\;\;={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})\,{\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ (S141)
$\qquad\quad\;\;={\mathcal{N}}\left(\mu=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},\sigma^{2}=(\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}})^{-1}\right),$ (S142)
where we define the hyperparameters $\mu_{o}$ and $\tau_{o}$ as follows:
$\mu_{o}=\frac{1}{J}\sum_{y_{\hat{j}}\in N_{J}({\bf{x}}_{o})}y_{\hat{j}}\qquad\text{and}\qquad\tau_{o}^{-1}=\frac{1}{J-1}\sum_{y_{\hat{j}}\in N_{J}({\bf{x}}_{o})}(y_{\hat{j}}-\mu_{o})^{2}.$
Here, $N_{J}({\bf{x}}_{o})$ denotes a set consisting of $J$ closest neighbors to ${\bf{x}}_{o}$. Specifically, we compute a kernel function $k({\bf{x}}_{o},{\mathcal{D}}\,\setminus\bar{{\mathcal{D}}})\in^{1\times(n-\bar{o})}$, where $\bar{{\mathcal{D}}}$ represents inliers. Then, we retrieve $\{{\bf{x}}_{1},\dots,{\bf{x}}_{J}\}$ corresponding with the $J$ largest $k({\bf{x}}_{o},{\mathcal{D}}\,\setminus\bar{{\mathcal{D}}})\in^{1\times(n-\bar{o})}$. We set $J=3$ for our experiments, ensuring each outlier has four neighbors. This posterior also maintains conjugacy, yielding a normally distributed posterior. For each identified outlier $y_{o}$, we define the prior for the latent variable $\bar{\mu}_{o}$ using the sample mean and variance of its neighbors. Intuitively, this prior reflects the assumption that a human expert’s correction for a particularly identified outlier will be close to its neighbors. This prior resembles a probabilistic model inspired by case-based reasoning. We then obtain the expected value ${\mathbb{E}}_{p(\bar{\mu}_{o}|\bar{y}_{o})}$ as
${\mathbb{E}}_{p(\bar{\mu}_{o}|\bar{y}_{o})}=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},$ (S143)
which equals the posterior mean. The algorithm to implement the informative mean prior is presented in Algorithm 1.
Algorithm 1 Expert-guided robust mean prior algorithm
1:  Input: ${\mathcal{D}}$, $\bar{{\mathcal{D}}}$, $\bar{{\bf{o}}}$, $\bar{{\bf{y}}}$, $\sigma^{2}_{\mathrm{corr.}}$
2:  $\bar{\mu}\leftarrow\frac{1}{|{\mathcal{D}}|}\sum_{j=1}^{|{\mathcal{D}}|}\,y_{j}$
3:  $\bar{\sigma}\leftarrow\sqrt{\frac{1}{|{\mathcal{D}}|-1}\sum_{j=1}^{|{\mathcal{D}}|}(y_{j}-\bar{\mu})^{2}}$
4:  for $o=1,\dots,\hat{o}$ do
5:     $\alpha_{o}\leftarrow|(\hat{y}_{o}-\bar{\mu})/\bar{\sigma}|$
6:     $\beta_{o}\leftarrow 0$
7:     $\mu_{o}\leftarrow\mu_{o}=\frac{1}{J}\sum_{y_{\hat{j}}\in N_{J}({\bf{x}}_{o})}y_{\hat{j}}$
8:     $\tau_{o}\leftarrow\frac{1}{J-1}\sum_{y_{\hat{j}}\in N_{J}({\bf{x}}_{o})}(y_{\hat{j}}-\mu_{o})^{2}$
9:     ${\mathbb{E}}_{p(\bar{\delta}_{o}|\bar{o}_{o})}[\bar{\delta}_{o}]\leftarrow\frac{\alpha_{o}+\bar{o}_{o}}{\alpha_{o}+\beta_{o}+1}$
10:     ${\mathbb{E}}_{p(\bar{\mu}_{o}|\bar{y}_{o})}[\bar{\mu}_{o}]\leftarrow\frac{\tau_{o}\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+\sigma^{-2}_{\mathrm{corr.}}}$
11:  end for
12:  $m({\bf{x}})\leftarrow\frac{1}{\hat{o}}\sum_{o}\,{\mathbb{E}}_{p(\bar{\delta}_{o}|\bar{o}_{o})}[\bar{\delta}_{o}]\;{\mathbb{E}}_{p(\bar{\mu}_{o}|\bar{y}_{o})}[\bar{\mu}_{o}]$
13:  Output: mean prior $m({\bf{x}})$
Appendix H Benchmark details
H.1 UCI regression datasets
Boston
The dataset consists of $n=506$ observations, each representing a suburban or town area in Boston. It encompasses $d=13$ features containing data like the average number of rooms in dwellings, pupil-teacher ratios, and per capita crime rates. We try to predict the median price of homes residents own (excluding rented properties). The dataset can be found at https://www.cs.toronto.edu/~delve/data/boston/bostonDetail.html.
Energy
The dataset describes the energy efficiency of buildings by correlating their heating and cooling load requirements with various building parameters. It consists of $n=768$ data samples, each characterised by $d=8$ distinct features, with the ultimate goal of predicting a single continuous response variable found in the last column. The dataset can be found at https://archive.ics.uci.edu/dataset/242/energy+efficiency.
Yacht
The dataset’s main focus is on predicting the residuary resistance of sailing yachts during their initial design phase, a critical aspect in evaluating a vessel’s performance and estimating the essential propulsive power required. This prediction relies on $d=6$ primary input parameters, which include the fundamental hull dimensions and boat velocity. The dataset contains $n=308$ observations. The dataset can be found at https://archive.ics.uci.edu/dataset/243/yacht+hydrodynamics.
Parkinsons
The Parkinsons Telemonitoring dataset contains $n=197$ data samples and $d=22$ input features derived from voice measures. The dataset comprises a collection of biomedical voice measurements from 31 individuals, including 23 diagnosed with Parkinson’s disease (PD). Each row represents one of 195 voice recordings, while each column corresponds to a specific vocal feature.. The dataset can be found at https://archive.ics.uci.edu/dataset/174/parkinsons.
H.2 High-throughput Bayesian Optimization problems
Hartmann 6D.
The widely used Hartmann benchmark function [38].
Lunar Lander.
The goal of this task is to find an optimal $12$-dimensional control policy that allows an autonomous lunar lander to consistently land without crashing. The final objective value we optimize is the reward obtained by the policy averaged over a set of 50 random landing terrains. For this task, we use the same controller setup used by [8].
Rover.
The rover trajectory optimization task introduced by [44] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a set of obstacles. We use the same obstacle set up as in [29].
Lasso DNA.
We optimize the $180-$dimensional DNA task from the LassoBench library [49] of benchmarks based on weighted LASSO regression.
H.3 Additional outlier contamination protocols on UCI regression datasets
Our outlier contamination protocols follow the same settings as [2]. We recall the asymmetric outlier case described in the main text in Section 7.
Asymmetric
We sample uniformly at random 10% of the training dataset input-output pairs $({\bf{x}}_{i},y_{i})$, and replace the $y_{i}$’s by asymmetric outliers, i.e., via subtraction of noise sampled from a uniform distribution $\mathcal{U}(3\bar{\sigma},9\bar{\sigma})$, with $\bar{\sigma}$ being the standard deviation of the original observations.
Next, Section 7.3 considers two additional kind of outliers:
Uniform
Instead of always subtracting noise as was the case for asymmetric outliers, the set of training data points selected for outlier contamination is divided in 2 subsets: half of the selected subset is contaminated by adding $z\sim U(3\sigma,9\sigma)$, while the other half is contaminated by subtracting $z\sim U(3\sigma,9\sigma)$.
Focused
In this outlier generation process, we randomly select and remove a subset of data points, which will be replaced by outliers. For these outliers, we deterministically choose their values in $\mathcal{X}$. To do so, we calculate the median value for each input data dimension $j$. However, we do not place the outliers at this median position directly. Instead, we replace the removed input values by $(m_{1}+\delta_{1},m_{2}+\delta_{2}\dots,m_{d}+\delta_{d})^{\top}$, where $m_{j}$ is the median in the $j$-th input data dimension, and $\delta_{j}=\alpha_{j}u$, where $\alpha_{j}$ is the median absolute deviation of the $j$-th data dimension times 0.1, and $u\sim U(0,1)$. Simultaneously, the outlier values on $\mathcal{Y}$ are obtained by subtracting three times the standard deviation of the median of the observations $M_{y}$. To not have the same value for every outlier position, we also add a small perturbation $\delta_{y}=\alpha_{y}u$, where $\alpha_{y}$ is the median absolute deviation of ${\bf{y}}$ times 0.1, and $u\sim U(0,1)$.
Appendix I Experimental Details
I.1 Hyperparameters
UCI regression
Optimizer ADAM
Learning rate $0.01$
Minibatch size $n$-data
Number of iterations for optimizing ELBO $50$
Proportion of test set $0.2$
$\epsilon$ $0.2$
$p$-outliers $0.1$
$n$-inducing points (SVGP) 100
Projection-dim (RCaGP) 5
Mean-prior $m$ (RCaGP and RCSVGP) $\frac{1}{n}\sum_{j}^{n}y_{j}$
Mean-prior $m$ (SVGP) 0
$\beta$ $1.0$
$\sigma^{2}_{\mathrm{corr.}}$ 1.0
High-throughput BO
$n_{0}$ (data initialization) 250
ADAM step size for query ${\bf{x}}$ 0.001
ADAM step size for RCaGP parameters 0.01
The number of expert corrections $\text{T}_{\text{iterations}}/20$
Minibatch size 250
$\sigma^{2}_{\mathrm{corr.}}$ $1.0$
$p-$outliers 0.25
Mean-prior $m$ (all models) $0.0$
Projection-dim (RCaGP) 25
$n$-inducing points (SVGP) 25
Table S3: Hyperparameter settings used for RCaGP. Most of them remain consistent across all the tasks.
I.2 Hardware
For the UCI regression experiments, all models—including our proposed method and the baselines—were executed on a compute cluster consisting of two machines: one equipped with dual AMD EPYC 7713 processors (64 cores each, 2.0 GHz), and another with dual Intel Xeon Gold 6148 processors (20 cores each, 2.4 GHz). The longest-running UCI task, based on the Energy dataset, required approximately 15 minutes to complete, including both our model and all baseline methods. Bayesian Optimization experiments were conducted on a cluster with four NVIDIA V100 GPUs, each with 32 GB of memory. Among these tasks, the Hartmann 6D benchmark was the fastest, completing in roughly 20 minutes, while the most computationally intensive task—DNA—required up to 20 hours of runtime.

### v2

@@
 1 Introduction
-Gaussian Processes (GPs) are a foundational tool in probabilistic machine learning, offering non-parametric modeling with principled uncertainty estimates [35].
+Gaussian Processes (GPs) are a foundational tool in probabilistic machine learning, offering non-parametric modeling with principled uncertainty estimates [Rasmussen2006].
 Their use spans diverse domains such as time series forecasting, spatial modeling, and regression on structured scientific data.
-One particularly impactful area where GPs have become integral is Bayesian Optimization (BO), a sample-efficient framework for optimizing expensive black-box functions, where the surrogate model’s predictive accuracy and uncertainty calibration directly influence decision-making [10].
+One particularly impactful area where GPs have become integral is Bayesian Optimization (BO), a sample-efficient framework for optimizing expensive black-box functions, where the surrogate model’s predictive accuracy and uncertainty calibration directly influence decision-making [garnett2023bayesian].
 Many scientific and industrial tasks are high-throughput optimization problems in which numerous experiments or simulations are performed, possibly in parallel.
-These can be effectively tackled by BO and include material design [14], drug discovery with extensive chemical space search [17], and fine-tuning gene expression in synthetic biology [5].
+These can be effectively tackled by BO and include material design [hastings2024interoperable], drug discovery with extensive chemical space search [hernandez2016distributed], and fine-tuning gene expression in synthetic biology [belanger2019biological].
 Yet, two major challenges limit the reliability of GPs in both regression and BO settings: scalability to large datasets and robustness to corrupted or outlier-contaminated observations.
-Sparse Variational Gaussian Processes (SVGPs, [16, 40, 41]) address scalability via inducing point approximations, yielding a low-rank approximation that allows large-scale regression and high-throughput BO.
+Sparse Variational Gaussian Processes (SVGPs, [hensman2013gaussian, titsias2009variational, titsias2014doubly]) address scalability via inducing point approximations, yielding a low-rank approximation that allows large-scale regression and high-throughput BO.
 Figure 1: Overview of our proposed RCaGP against concurrent baselines on a 1D example.
@@
 Overconfident predictions occur in regions where the inducing points are far from observed data.
-Accurate uncertainty quantification is essential for sequentially identifying the optimal data points in BO [44].
+Accurate uncertainty quantification is essential for sequentially identifying the optimal data points in BO [wang2018batched].
 Contributions.
-To address these challenges, we propose a novel approach that combines robust-conjugate SVGP [2] with computation-aware GP [46].
-This method, referred to as the Robust Computation-Aware GP (RCaGP), jointly tackles outlier robustness and approximation-induced uncertainty—two challenges that are deeply intertwined in sparse GP models.
+To address these challenges, we propose a novel approach that combines robust-conjugate SVGP [altamirano2023robust] with computation-aware GP [wenger2024computation].
+This method, referred to as the Robust Computation-Aware GP (RCaGP), jointly tackles outlier robustness and approximation-induced uncertainty, two challenges that are deeply intertwined in sparse GP models.
 Computational approximations such as low-rank matrix factorizations can exacerbate the effects of outliers, and mitigating one without addressing the other leads to suboptimal performance.
@@
   1. 1.
-    (Methodological) introduces a novel GP model (RCaGP) that jointly addresses approximation-awareness and outlier robustness through a principled probabilistic framework—going beyond previous work that tackles these issues in isolation.
+    (Methodological) introduces a novel GP model (RCaGP) that jointly addresses approximation-awareness and outlier robustness through a principled probabilistic framework, going beyond previous work that tackles these issues in isolation.
   2. 2.
@@
 The Gaussian noise assumption in GPs often leads to misspecification, e.g., due to outliers.
-To address this problem, recent work proposed robust conjugate GP (RCGP), which employs a robust loss function through generalized Bayesian inference [2].
+To address this problem, recent work proposed robust conjugate GP (RCGP), which employs a robust loss function through generalized Bayesian inference [altamirano2023robust].
 Consider a noisy dataset ${\mathcal{D}}_{n}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}=({\bf{X}},{\bf{y}})$.
 We specify a GP prior for the finite function evaluations ${\bf{f}}=[f({\bf{x}}_{1}),\dots,f({\bf{x}}_{n})]$: $p({\bf{f}})={\mathcal{N}}({\bf{f}};\!{\bf{m}},\!{\bf{K}})$, with ${\bf{m}}=[m({\bf{x}}_{1}),\dots,m({\bf{x}}_{n})]^{\top}$ and ${\bf{K}}=[k({\bf{x}}_{j},\!{\bf{x}}_{l})]_{1\leq j,l\leq n}$.
-Both $m:{\mathcal{X}}\rightarrow$ and $k:{\mathcal{X}}\times{\mathcal{X}}\rightarrow$ denote the mean and kernel functions, encoding prior beliefs about the function’s structure and smoothness [35].
+Both $m:{\mathcal{X}}\rightarrow$ and $k:{\mathcal{X}}\times{\mathcal{X}}\rightarrow$ denote the mean and kernel functions, encoding prior beliefs about the function’s structure and smoothness [Rasmussen2006].
 Conditioning on observations, the posterior predictive distribution is Gaussian with mean $\mu_{\ast}({\bf{x}})$ and variance $k_{\ast}({\bf{x}},\!{\bf{x}})$.
@@
 Nevertheless, RCGP inherits the cubic inference complexity of vanilla GPs, prohibiting its usage on large-scale datasets like those encountered in high-throughput BO.
-To remedy this issue, RCSVGP were introduced, enabling outlier-robust regression on large-scale data [2].
+To remedy this issue, RCSVGP were introduced, enabling outlier-robust regression on large-scale data [altamirano2023robust].
 This model introduces inducing points ${\bf{u}}=\{{\bf{u}}_{l}\}_{l=1}^{r}$ for $r\ll n$, with inducing locations ${\bf{z}}=\{{\bf{z}}_{l}\}_{l=1}^{r}$ such that ${\bf{u}}_{l}=f({\bf{z}}_{l})$.
@@
 Computation-aware GP.
-[45, 46] proposed computation-aware GP (CaGP) to address the SVGP overconfidence.
+[wenger2022posterior, wenger2024computation] proposed computation-aware GP (CaGP) to address the SVGP overconfidence.
 CaGP mitigates this issue by ensuring its posterior variance exceeds the vanilla GP variance.
@@
 ${\boldsymbol{\Sigma}}_{i}={\bf{K}}^{-1}-{\bf{C}}_{i},$ (5)
-where ${\bf{C}}_{i}={\bf{S}}_{i}({\bf{S}}_{i}^{\top}\hat{{\bf{K}}}^{-1}{\bf{S}}_{i}){\bf{S}}_{i}\approx{\bf{K}}^{-1}$ is a low-rank approximation and ${\bf{S}}_{i}\in^{n\times i}$ is the action matrix associated with the chosen approximation technique.
+where ${\bf{C}}_{i}={\bf{S}}_{i}({\bf{S}}_{i}^{\top}\hat{{\bf{K}}}^{-1}{\bf{S}}_{i}){\bf{S}}_{i}^{\top}\approx{\bf{K}}^{-1}$ is a low-rank approximation and ${\bf{S}}_{i}\in^{n\times i}$ is the action matrix associated with the chosen approximation technique.
 In the case of inducing points, ${\bf{S}}_{i}=k({\bf{X}},{\bf{z}}_{i})$.
@@
 As $i\rightarrow n$, CaGP recovers both the predictive mean and variance of the exact GP.
-Bayesian optimization.
-BO is a sequential algorithm designed to solve the problem ${\mathrm{argmax}}_{{\bf{x}}\in{\mathcal{X}}}f({\bf{x}})$ for ${\mathcal{X}}$ a compact domain, and $f$ a black-box function, meaning only noisy evaluations are accessible [10].
-BO involves a probabilistic surrogate $p(f|{\mathcal{D}})$ and an acquisition function (AF) $\alpha({\bf{x}}):=\int u({\bf{x}};f)\,p(f|{\mathcal{D}})\,\mathrm{d}f$, where $u$ measures the usefulness of evaluating $f({\bf{x}})$.
-GPs are usually employed as the surrogate, and popular AFs include, for instance, expected improvement, knowledge gradient, and entropy search [10].
-Importantly, these AFs can be maximized in a batch setting, yielding $q$ candidates every iteration $t,{\bf{X}}_{t}^{q}={\mathrm{argmax}}_{{\bf{X}}\in{\mathcal{X}}^{q}}\alpha({\bf{X}})$ [48].
 3 Robust and Computation-Aware Gaussian Processes
@@
 3.1 Robust Computation-aware Inference
-Inspired by CaGP, we apply a probabilistic treatment to the representer weights $\hat{{\bf{v}}}=({\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}\!-\!{\bf{m}}_{\bf{w}})$ of RCGP in Equation 1.
+Inspired by CaGP, we apply a probabilistic treatment to the representer weights $\hat{{\bf{v}}}=({\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}\!-\!{\bf{m}}_{\bf{w}})$ of RCGP in Equation˜1.
 This enables us to quantify uncertainty in approximating $\hat{{\bf{v}}}$—an aspect most approximate GPs ignore.
@@
 We begin by placing a prior on $\hat{{\bf{v}}}$, i.e., $\hat{{\bf{v}}}\sim{\mathcal{N}}(\tilde{{\bf{v}}}_{0}\!=\!\mathbf{0},\,\tilde{{\boldsymbol{\Sigma}}}_{0}\!=\!\tilde{{\bf{K}}}^{-1})$, where $\tilde{{\bf{K}}}={\bf{K}}\!+\!{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}}$.
-We then update this belief via the linear Gaussian identity [45], yielding a Gaussian posterior with mean $\tilde{{\bf{v}}}_{i}$ and variance $\tilde{{\boldsymbol{\Sigma}}}_{i}$:
+We then update this belief via the linear Gaussian identity [wenger2022posterior], yielding a Gaussian posterior with mean $\tilde{{\bf{v}}}_{i}$ and variance $\tilde{{\boldsymbol{\Sigma}}}_{i}$:
 $\tilde{{\bf{v}}}_{i}=\tilde{{\bf{C}}}_{i}\,({\bf{y}}-{\bf{m}}_{\bf{w}}),$ (8)
@@
 where $\tilde{{\bf{C}}}_{i}\!=\!{\bf{S}}_{i}({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})^{-1}{\bf{S}}_{i}^{\top}$ and ${\bf{S}}_{i}\!\in\!^{n\times i}$ are low-rank approximations of $\tilde{{\bf{K}}}^{-1}$ and the actions, respectively.
-The updates can be done sequentially [45, Algorithm 1] or in batches [46, Algorithm S2].
-Recent work generalized the action matrix ${\bf{S}}_{i}\in^{n\times i}$ by treating it as a variable that can be optimized during model selection [46].
+The updates can be done sequentially [wenger2022posterior, Algorithm 1] or in batches [wenger2024computation, Algorithm S2].
+Recent work generalized the action matrix ${\bf{S}}_{i}\in^{n\times i}$ by treating it as a variable that can be optimized during model selection [wenger2024computation].
 Given a new data point ${\bf{x}}\in{\mathcal{X}}$, the predictive distribution of $f({\bf{x}})$ follows a normal distribution with mean $\hat{\mu}_{i}({\bf{x}})$ and variance $\hat{k}_{i}({\bf{x}},{\bf{x}})$:
@@
 $\hat{k}_{i}({\bf{x}},{\bf{x}})=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}\tilde{{\bf{K}}}_{i}{\bf{k}}_{\bf{x}}+{\bf{k}}_{\bf{x}}^{\top}\tilde{{\boldsymbol{\Sigma}}}_{i}{\bf{k}}_{\bf{x}}=k({\bf{x}},{\bf{x}})-{\bf{k}}_{\bf{x}}^{\top}\tilde{{\bf{C}}}_{i}{\bf{k}}_{\bf{x}}.$ (11)
-In Equation 11, RCaGP incorporates the computational uncertainty ${\bf{k}}_{\bf{x}}^{\top}\,\tilde{{\boldsymbol{\Sigma}}}_{i}\,{\bf{k}}_{\bf{x}}$, corresponding to the variance of the representer weights $\tilde{{\boldsymbol{\Sigma}}}_{i}$, resulting in a combined uncertainty $\hat{k}_{i}({\bf{x}},{\bf{x}})$.
+In Equation˜11, RCaGP incorporates the computational uncertainty ${\bf{k}}_{\bf{x}}^{\top}\,\tilde{{\boldsymbol{\Sigma}}}_{i}\,{\bf{k}}_{\bf{x}}$, corresponding to the variance of the representer weights $\tilde{{\boldsymbol{\Sigma}}}_{i}$, resulting in a combined uncertainty $\hat{k}_{i}({\bf{x}},{\bf{x}})$.
 This computational uncertainty guarantees that the predictive posterior variance of RCaGP is larger than RCGP, preventing RCaGP from overconfident predictions.
 Such behavior is referred to as a conservative uncertainty estimate.
-The computational complexity of RCaGP can be found in Appendix B.
-The robustness of RCaGP hinges on the weight function $w$, which we choose to incorporate into the representer weights $\tilde{{\bf{v}}}$. To that end, we follow [2] and let:
+The computational complexity of RCaGP can be found in Appendix˜B.
+The robustness of RCaGP hinges on the weight function $w$, which we choose to incorporate into the representer weights $\tilde{{\bf{v}}}$. To that end, we follow [altamirano2023robust] and let:
 $w({\bf{x}},y)=\beta\left(1+(y-m({\bf{x}}))^{2}/c^{2}\right)^{-1/2},$ (12)
 for $\beta$ and $c>0$ the learning rate and soft threshold.
-The weight function assigns smaller values to observations $y$ that deviate significantly from the mean prior $m({\bf{x}})$, treating them as potential outliers.
-Lower weights reduce the influence of such points during inference, leading to a more robust posterior.
-Additionally, the weight enters the noise term ${\bf{J}}_{\bf{w}}=\mathrm{diag}([{\sigma_{\mathrm{noise}}^{2}}/2{\bf{w}}^{-2}])$, effectively inflating the noise for outliers and further limiting their impact.
-$w({\bf{x}},y)$ depends critically on the mean prior: for an outlier $\hat{y}$, if $m(\hat{{\bf{x}}})$ is close to $\hat{y}$, the weight increases.
-However, choosing an informative mean prior is difficult without domain knowledge.
+[…]
 3.2 Model hyperparameters optimization
 We use the evidence lower bound (ELBO) as a loss function to optimize the kernel hyperparameters ${\boldsymbol{\theta}}\!\in\!^{p}$.
-Following [46], this enables RCaGP to scale to large datasets while avoiding overconfidence.
+Following [wenger2024computation], this enables RCaGP to scale to large datasets while avoiding overconfidence.
 The variational family is defined using the RCaGP posterior $q_{i}({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})$, i.e., $\{q_{i}({\bf{f}})={\mathcal{N}}(\hat{\mu}_{i}({\bf{X}}),\hat{k}_{i}({\bf{X}},\!{\bf{X}}))|{\bf{S}}_{i}\!\in\!^{n\times i}\}$, parameterized by action ${\bf{S}}_{i}$.
-We replace the evidence term $\log p({\bf{y}}|{\boldsymbol{\theta}})$ with the robust loss $\log\Psi^{w}({\bf{y}},\!{\bf{f}})$ from [2].
+We replace the evidence term $\log p({\bf{y}}|{\boldsymbol{\theta}})$ with the robust loss $\log\Psi^{w}({\bf{y}},\!{\bf{f}})$ from [altamirano2023robust].
 We then formulate ELBO as
@@
 This loss learns the hyperparameters as if maximizing ${\mathbb{E}}_{p({\bf{f}}|{\bf{y}},{\boldsymbol{\theta}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ while minimizing the computational uncertainty associated with the approximation error.
-The derivation of the expected loss ${\mathbb{E}}_{q({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ follows [2], while the derivation of the KL term ${\mathrm{KL}}[q({\bf{f}})\|p({\bf{f}})]$ draws an analogy to CaGP.
-For the closed form and detailed derivation of ELBO, we refer the readers to Proposition C.1.
+The derivation of the expected loss ${\mathbb{E}}_{q({\bf{f}})}[\log\Psi^{w}({\bf{y}},{\bf{f}})]$ follows [altamirano2023robust], while the derivation of the KL term ${\mathrm{KL}}[q({\bf{f}})\|p({\bf{f}})]$ draws an analogy to CaGP.
+For the closed form and detailed derivation of ELBO, we refer the readers to Proposition˜C.1.
 We obtain the optimal hyperparameters $\boldsymbol{\theta}$$^{\ast}$ by minimizing the negative ELBO.
 3.3 Joint model parameters and design selection strategy
-To fully take advantage of RCaGP in BO, we leverage the expected lower bound utility (EULBO) framework introduced by [28].
+To fully take advantage of RCaGP in BO, we leverage the expected lower bound utility (EULBO) framework introduced by [maus2024approximation].
 At acquisition time, instead of optimizing the surrogate’s inducing points, variational parameters, and hyperparameters, only then to find a design maximizing the acquisition function, a global criterion is maximized in an end-to-end manner:
@@
 This reformulation replaces the commonly used ReLU function with a softplus function, guaranteeing the utility function remains strictly positive whenever $f({\bf{x}})\geq y^{\ast}$.
-Moreover, EULBO can be extended to support batch BO using Monte Carlo batch mode [4, 47].
+Moreover, EULBO can be extended to support batch BO using Monte Carlo batch mode [balandat2020botorch, wilson2018maximizing].
 Given a set of candidates ${\bf{X}}=\{{\bf{x}}_{1},\dots,{\bf{x}}_{q}\}$, the expected utility function corresponding to the $q$-improvement utility is:
@@
 ${\bf{x}}_{1}^{\ast},\dots,{\bf{x}}_{q}^{\ast},{\bf{S}}_{i}^{\ast},{\boldsymbol{\theta}}_{t}^{\ast}=\underset{{\bf{x}}_{1},\dots,{\bf{x}}_{q},{\bf{S}},{\boldsymbol{\theta}}}{{\mathrm{argmax}}}\;\ell_{\mathrm{EULBO}}^{\mathrm{RCaGP}}.$ (17)
-For efficiency, we impose the sparse block structure on the action matrices ${\bf{S}}_{i}\in^{n\times i}$, following [46].
+For efficiency, we impose the sparse block structure on the action matrices ${\bf{S}}_{i}\in^{n\times i}$, following [wenger2024computation].
 Specifically, we enforce each block to be column vector ${\bf{s}}_{j}\in^{k\times 1}$, with $k=n/i$ entries, so that the number of trainable parameters is $k\times i=n$ (training data size).
 It is worth noticing that unlike EULBO with SVGP, which maximizes variational parameters and inducing locations, our approach directly optimizes the action matrices ${\bf{S}}_{i}$.
-Maximizing the ELBO independently of the posterior-expected utility function can result in suboptimal data acquisition decisions, as the ELBO is primarily designed to model observed data [27].
-In contrast, EULBO considers how the surrogate performs when selecting the next query, effectively guiding the solution of ELBO optimization towards high utility regions [28].
+Maximizing the ELBO independently of the posterior-expected utility function can result in suboptimal data acquisition decisions, as the ELBO is primarily designed to model observed data [matthews2016sparse].
+In contrast, EULBO considers how the surrogate performs when selecting the next query, effectively guiding the solution of ELBO optimization towards high utility regions [maus2024approximation].
 Our RCaGP model further balances such aggressive behavior by incorporating the combined uncertainty, preventing overconfidence during candidate query selection.
@@
 This section presents a theoretical analysis of RCaGP.
-Proposition 4.1 establishes its robustness to outliers, while Proposition 4.2 links its uncertainty estimates to worst-case error, ensuring conservative predictions, thus addressing the challenges outlined in Section 1.
-We demonstrate the robustness of RCaGP through the posterior influence function (PIF), similarly to [2, Proposition 3.2].
-The PIF is a criterion proving robustness to misspecification in observation error [11, 26].
+Proposition˜4.1 establishes its robustness to outliers, while Proposition˜4.2 links its uncertainty estimates to worst-case error, ensuring conservative predictions, thus addressing the challenges outlined in Section˜1.
+We demonstrate the robustness of RCaGP through the posterior influence function (PIF), similarly to [altamirano2023robust, Proposition 3.2].
+The PIF is a criterion proving robustness to misspecification in observation error [ghosh2016robust, matsubara2022robust].
 For this purpose, we define the dataset ${\mathcal{D}}=\{({\bf{x}}_{j},y_{j})\}_{j=1}^{n}$ and the corresponding contaminated dataset ${\mathcal{D}}_{m}^{c}=({\mathcal{D}}\setminus({\bf{x}}_{m},y_{m}))\cup({\bf{x}}_{m},y_{m}^{c})$, indexed by $m\in\{1,\dots,n\}$.
@@
 Here, we choose KL-divergence since it provides a closed-form expression for multivariate Gaussian.
-A posterior is robust if $\sup_{y\in\mathcal{Y}}|\mathrm{PIF}(y_{m}^{c},{\mathcal{D}})|<\infty$ [2].
+A posterior is robust if $\sup_{y\in\mathcal{Y}}|\mathrm{PIF}(y_{m}^{c},{\mathcal{D}})|<\infty$ [altamirano2023robust].
 We then show that RCaGP’s PIF is bounded through the following proposition:
@@
 The constraint $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$ ensures no observation has infinite weight, and $\sup_{{\bf{x}},y}y,w^{2}({\bf{x}},y)<\infty$ guarantees that $w$ down-weights observations at least at rate $1/y$.
-The proof sketch shows that each term in Equation 18 is bounded.
+The proof sketch shows that each term in Equation˜18 is bounded.
 We leverage the pseudo-inverse property of positive semi-definite matrices and matrix norm bounds on the low-rank approximation to relate RCGP and CaGP.
-The full proof is in Appendix D.
-Crucially, the PIF of CaGP is unbounded, indicating a lack of robustness to outliers (see Section F.1).
-Next, we show that RCaGP’s uncertainty captures the worst-case error over all latent functions, paralleling [45, Theorem 2].
+The full proof is in Appendix˜D.
+Crucially, the PIF of CaGP is unbounded, indicating a lack of robustness to outliers (see Section˜F.1).
+Next, we show that RCaGP’s uncertainty captures the worst-case error over all latent functions, paralleling [wenger2022posterior, Theorem 2].
 We assume the difference between the latent function $g$ and the shrinkage term $m_{w}$ lies in an RKHS.
+This holds when $\|\nabla_{y}\log w^{2}({\bf{x}},y)\|^{2}_{{\mathcal{H}}_{k}}<\infty$.
+In the case of our robust weight function, the correction term is a nonlinear rational function.
+Generally, such functions do not lie in the RKHS associated with common kernels (e.g., RBF, Matérn), unless specific and uncommon conditions are met—such as the RKHS being closed under composition with the given nonlinearity.
+A practical approach to ensure it belongs to the RKHS is to project the correction term onto the RKHS.
+Alternatively, one could use or design a kernel whose RKHS explicitly accommodates this class of nonlinear transformations.
 Under this assumption, we establish the link between RCaGP’s uncertainty estimates and the worst-case error through the following proposition:
@@
 where $\mu_{\ast}^{g}(\cdot)=k(\cdot,{\bf{X}})\tilde{{\bf{K}}}^{-1}h({\bf{X}})$ is the RCGP’s posterior and $\hat{\mu}_{i}^{g}(\cdot)=k(\cdot,{\bf{X}})\tilde{{\bf{C}}}_{i}(g({\bf{X}})-m_{w}({\bf{X}}))$ RCaGP’s posterior mean for a function $h\in{\mathcal{H}}_{k^{w}}=g({\bf{x}})-m_{w}({\bf{x}})$ with a latent function $g:{\mathcal{X}}\rightarrow$ and the shrinkage function $m_{w}:{\mathcal{X}}\times{\mathcal{Y}}\rightarrow$.
-The first equation in Proposition 4.2 shows that RCaGP’s uncertainty $\hat{k}_{i}(.,.)$ captures the worst-case error between the latent function $g$ and the posterior mean $\hat{\mu}_{i}(.)$.
+The first equation in Proposition˜4.2 shows that RCaGP’s uncertainty $\hat{k}_{i}(.,.)$ captures the worst-case error between the latent function $g$ and the posterior mean $\hat{\mu}_{i}(.)$.
+Instead of pretending the model is fully correct, we allow a structured form of misspecification, namely, the part captured by ${\bf{m}}_{w}$.
+Proposition˜4.2 then guarantees that any remaining discrepancy is still bounded by our variance.
+Even when classical Bayesian inference may fail to offer meaningful uncertainty due to model misspecification or outliers, RCaGP maintains a rigorous worst-case interpretation of its predictive variance, so long as the residual is sufficiently regular.
+This mirrors the CaGP assumption in spirit: just as CaGP analyzes worst-case error over functions in the RKHS, RCaGP does the same for the residual, which can be viewed as the portion of the signal that the robust update leaves uncorrected.
 The second shows that the computational uncertainty $\hat{\sigma}_{i}^{2}$ captures the worst-case error between the RCGP and RCaGP posterior means, induced by approximation actions such as inducing points.
-This added uncertainty prevents overconfidence when modeling the latent function.
-The proof follows [45], leveraging [Proposition 3.9] of [20] and the RKHS associated with RCGP.
-Full details are in Appendix E.
-Lastly, the convergence of RCaGP in terms of mean function in RKHS norm is established in Section F.2.
+The proof follows [wenger2022posterior], leveraging [kanagawa2018gaussian, Proposition 3.9] and the RKHS associated with RCGP.
+Full details are in Appendix˜E.
+Lastly, the convergence of RCaGP in terms of mean function in RKHS norm is established in Section˜F.2.
 5 Expert-guided robust mean prior definition
-Our proposed RCaGP enforces outlier robustness through the weight function $w$, which—as discussed in Section 3.1 and Proposition 4.2—critically depends on the prior mean $m$.
+Our proposed RCaGP enforces outlier robustness through the weight function $w$, which—as discussed in Section 3.1 and Proposition˜4.2—critically depends on the prior mean $m$.
 Motivated by these insights, we propose defining a robust and informative mean prior using domain expert feedback.
@@
 In practice, we average over all identified outliers and use only posterior expectations, prioritizing computational efficiency over full Bayesian inference.
-The full generative process and toy illustration appear in Figure 2 (left and middle).
-Additional inference details are provided in Appendix G.
+The full generative process and toy illustration appear in Figure˜2 (left and middle).
+Additional inference details are provided in Appendix˜G.
 Figure 2: Expert-guided robust mean prior.
@@
 Outlier-robust Gaussian processes fall into three main categories.
-The first replaces Gaussian noise with heavy-tailed alternatives—e.g., Student-$t$[19, 34], Huber[1], Laplace [22], data-dependent [12], or mixture models [7, 24].
+The first replaces Gaussian noise with heavy-tailed alternatives—e.g., Student-tt[jylanki2011robust, ranjan2016robust], Huber[algikar2023robust], Laplace [kuss2006gaussian], data-dependent [goldberg1997regression], or mixture models [daemi2019identification, lu2023robust].
 These approaches break conjugacy and require approximate inference.
-The second class removes suspected outliers before applying standard inference [23, 31, 3], but this is often unreliable and computationally expensive in high dimensions.
-The third approach, RCGP [2], builds on generalized Bayesian inference [6, 18, 21] to down-weight outliers without discarding data.
+The second class removes suspected outliers before applying standard inference [li2021robust, park2022robust, andrade2023robust], but this is often unreliable and computationally expensive in high dimensions.
+The third approach, RCGP [altamirano2023robust], builds on generalized Bayesian inference [bissiri2016general, jewson2018principles, knoblauch2022optimization] to down-weight outliers without discarding data.
 We adopt RCGP as the foundation for RCaGP due to its strong performance in regression with outliers, its preserved conjugacy, and greater computational efficiency—enabling more accurate posterior inference and better predictive performance.
-SVGPs are the standard approach for scaling Gaussian processes to large-scale Bayesian optimization [13, 37, 43, 42].
-[30] improved inducing point selection by optimizing their weighted Gram matrix via a determinantal point process.
-[28] formulated acquisition as expected utility, enabling joint optimization of queries, variational parameters, hyperparameters, and inducing points in an end-to-end fashion.
-[25] proposed an online conditioning strategy that updates SVGPs with new data without reoptimizing variational parameters, enabling look-ahead acquisition in BO.
-Computation-aware Gaussian processes are a novel class of GPs that accounts for computational uncertainty, due to approximations in matrix-vector multiplications [45].
-These methods fall under the domain of probabilistic numerics, which quantify additional uncertainties arising from approximation errors [32, 39, 33].
-In that respect, [46] introduced a batch update algorithm for CaGP and proposed an ELBO objective to facilitate model selection.
-[15] extended CaGPs by proposing a calibrated probabilistic linear solver, resulting in a calibrated computation-aware GP model.
+SVGPs are the standard approach for scaling Gaussian processes to large-scale Bayesian optimization [griffiths2020constrained, stanton2022accelerating, vakili2021scalable, tripp2020sample].
+[moss2023inducing] improved inducing point selection by optimizing their weighted Gram matrix via a determinantal point process.
+[maus2024approximation] formulated acquisition as expected utility, enabling joint optimization of queries, variational parameters, hyperparameters, and inducing points in an end-to-end fashion.
+[maddox2021conditioning] proposed an online conditioning strategy that updates SVGPs with new data without reoptimizing variational parameters, enabling look-ahead acquisition in BO.
+Computation-aware Gaussian processes are a novel class of GPs that accounts for computational uncertainty, due to approximations in matrix-vector multiplications [wenger2022posterior].
+These methods fall under the domain of probabilistic numerics, which quantify additional uncertainties arising from approximation errors [pfortner2022physics, tatzel2023accelerating, pfortner2024computation].
+In that respect, [wenger2024computation] introduced a batch update algorithm for CaGP and proposed an ELBO objective to facilitate model selection.
+[hegde2024calibrated] extended CaGPs by proposing a calibrated probabilistic linear solver, resulting in a calibrated computation-aware GP model.
 7 Experimental results
-We evaluate the performance of our proposed RCaGP against concurrent baselines on a range of regression datasets, followed by several high-throughput Bayesian Optimization tasks.
+We evaluate the performance of our proposed RCaGP against concurrent baselines on a range of regression datasets, including a real-world dataset, followed by several high-throughput Bayesian Optimization tasks.
 We conclude by presenting insights gained after carrying out dedicated ablation studies.
+Our implementation is available at https://github.com/MarshalArijona/RCaGP.
 GP baselines.
-We compare RCaGP against RCSVGP [2], SVGP [16], and CaGP [46].
-Finally, we set the number of actions for RCaGP and CaGP, and the number of inducing points for SVGP and RCSVGP, to $25$.
-All baselines are equipped with a Matern $5/2$ kernel and were implemented using GPyTorch [9].
-Model hyperparameters are optimized using the ELBO (Equation 13).
+We compare RCaGP against several baseline methods: RCSVGP [altamirano2023robust], SVGP [hensman2013gaussian], and CaGP [wenger2024computation].
+For the UCI regression benchmarks, we further include comparisons with SVGP models using a Student-$t$ likelihood [jylanki2011robust] and SVGP with relevance pursuit (RRP), adapted from [ament2024robust] and formally derived in Section F.3.
+We set the number of actions (for RCaGP and CaGP) or inducing points (for SVGPs and RCSVGP) to 25.
+All baselines employ a Matérn-$5/2$ kernel and are implemented in GPyTorch [gardner2018gpytorch].
+Model hyperparameters are optimized by maximizing the ELBO (Equation 13).
 Outlier contamination protocol.
-In our experiments, we follow the settings described by [2].
+In our experiments, we follow the settings described by [altamirano2023robust].
 Unless stated otherwise in dedicated ablation studies, for regression datasets, we sample uniformly at random 10% of the training dataset input-output pairs $({\bf{x}}_{i},y_{i})$, and replace the $y_{i}$’s by asymmetric outliers, i.e., via subtraction of noise sampled from a uniform distribution $\mathcal{U}(3\bar{\sigma},9\bar{\sigma})$, with $\bar{\sigma}$ being the standard deviation of the original observations.
@@
 Bayesian Optimization settings.
-As described during Section 3.3, we integrate the baselines into EULBO, building on the implementation of [28].
+As described during Section 3.3, we integrate the baselines into EULBO, building on the implementation of [maus2024approximation].
 We initialize the optimization process for all baselines with $250$ data points sampled uniformly across the search space.
 All surrogates employ expected improvement (EI) as the acquisition function.
-Standard BO is applied across all tasks, while trust-region BO (TuRBO [8]) is additionally employed for high-dimensional tasks.
+Standard BO is applied across all tasks, while trust-region BO (TuRBO [eriksson2019scalable]) is additionally employed for high-dimensional tasks.
 7.1 Regression on UCI datasets
 We evaluate GP regression on four UCI datasets with asymmetric outliers (details in Section H.1); results are shown in the left part of Table 1.
-RCaGP consistently delivers the best performance, combining strong predictive accuracy and well-calibrated uncertainty with acceptable runtimes.
+RCaGP delivers the overall best performance, combining strong predictive accuracy and well-calibrated uncertainty with acceptable runtimes, and being only surpassed in the Yacht dataset by RRP, a much slower baseline.
 Notably, it significantly outperforms both CaGP and RCSVGP in this setting, indicating that the integration of robustness and approximation-awareness is effective.
-To test robustness across contamination types, we further consider two additional outlier scenarios, uniform and focused outliers, inspired by [2] (definitions in Section H.3).
+To test robustness across contamination types, we further consider two additional outlier scenarios, uniform and focused outliers, inspired by [altamirano2023robust] (definitions in Section H.3).
 The corresponding results are shown in Table S1 and Table S2.
@@
 These findings support its robustness across contamination types.
-More importantly, because outlier presence is not always known in advance, we also evaluate performance under clean conditions (right side of Table 1).
-RCaGP remains competitive and often superior—even without injected noise.
-One exception is the Energy dataset, where CaGP marginally outperforms RCaGP.
+More importantly, because outlier presence is not always known in advance, we also evaluate performance in the absence of outliers (right side of Table 1).
+RCaGP remains superior, except for the Energy dataset, where CaGP marginally outperforms RCaGP.
 This suggests that even in the absence of outliers, the weight function $w$ used by RCaGP (Equation 12) proves useful.
 An alternative explanation would be that these datasets already contain outliers from the start.
-Overall, these results demonstrate that RCaGP is more than the sum of its parts: the combination of RCGP and CaGP is not merely additive, but synergistic, consistently outperforming either component alone—both in outlier-rich and clean regimes.
+Overall, these results demonstrate that RCaGP is more than the sum of its parts: the combination of RCGP and CaGP is not merely additive, but synergistic, consistently outperforming either component alone, both in outlier-rich and clean regimes.
 Table 1: UCI Regression datasets results with asymmetric outliers and without outliers.
@@
 clock-time 2.450 ± 0.110 3.790 ± 0.130 0.870 ± 0.060 1.250 ± 0.090 2.580 ± 0.150 3.780 ± 0.190 0.910 ± 0.070 0.920 ± 0.020
+Student-t MAE 0.962 ± 0.037 1.015 ± 0.028 0.907 ± 0.061 1.025 ± 0.079 0.896 ± 0.041 0.970 ± 0.025 0.815 ± 0.071 0.965 ± 0.074
+NLL 1.619 ± 0.039 1.606 ± 0.024 1.564 ± 0.067 1.658 ± 0.085 1.593 ± 0.063 1.572 ± 0.026 1.559 ± 0.100 1.627 ± 0.092
+clock-time 1.210 ± 0.240 1.800 ± 0.400 1.350 ± 0.260 0.990 ± 0.090 1.330 ± 0.050 1.330 ± 0.080 0.890 ± 0.150 0.730 ± 0.020
+RRP MAE 0.550 ± 0.050 0.597 ± 0.112 0.356 ± 0.051 0.644 ± 0.106 0.815 ± 0.103 0.775 ± 0.118 0.708 ± 0.035 1.052 ± 0.100
+NLL 1.147 ± 0.050 1.240 ± 0.117 0.896 ± 0.030 1.306 ± 0.121 1.527 ± 0.194 1.547 ± 0.214 1.179 ± 0.045 / 1.884 ± 0.179
+clock-time 5.120 ± 0.100 3.840 ± 0.090 7.200 ± 0.160 3.010 ± 0.080 5.260 ± 0.120 3.730 ± 0.100 6.980 ± 0.110 3.010 ± 0.060
+7.2 Twitter Flash Crash
+We illustrate the real-world applicability of RCaGP by evaluating its performance and other baselines on the Dow Jones Industrial Average (DJIA) index for April 17, 2013.
+On this day, the Associated Press Twitter account was hacked, posting a false report of an explosion at the White House and an injury to the U.S.
+President.
+The resulting panic triggered a sudden market sell-off, followed by a rapid rebound, creating a brief interval during which the DJIA did not reflect the true economic state of the U.S.
+stock market.
+This scenario provides a unique stress test for evaluating model robustness under sudden, anomalous market conditions.
+Observations during this period can reasonably be treated as outliers.
+Figure˜3 (left) shows the DJIA index alongside the predictive means of RCaGP and other baselines.
+The predictive means of SVGP, RCSVGP, and CaGP exhibit noticeable deviations before or after 1 PM, while RCaGP yields a more reliable estimate, with only minor discrepancies around 2 PM.
+These results suggest that insufficient robustness or computational uncertainty in the baseline models contributes to their degraded performance under sudden market shocks.
+In the right panel of Figure˜3, we plot the weight functions of RCaGP and RCSVGP.
+Although their weight functions exhibit similar profiles, they yield markedly different mean predictions.
+This supports our claim that robustness alone, without an appropriate treatment of computational uncertainty, can degrade the performance of approximate Gaussian process models.
+Figure 3: Left: Predictive mean of various approximate GP models on the DJIA index with a constant mean prior.
+Outliers affect the baselines, whereas RCaGP maintains robustness.
+Right: learned weight functions $w({\bf{x}},y)$ for RCaGP and RCSVGP, illustrating their influence on the regression.
 7.3 High-Throughput Bayesian Optimization
 We consider 4 tasks: Hartmann6D, Lunar12D, Rover60D and Lasso-DNA180D (Section H.2).
-Figure 3 reports the best value found during BO, averaged over 20 trials, with $\pm$1 standard deviation.
+Figure˜4 reports the best value found during BO, averaged over 20 trials, with $\pm$1 standard deviation.
 To ensure fair comparisons despite outlier contamination, we report the uncontaminated mean best value.
@@
 SVGP and RCSVGP display the same performance.
-The bottom row (columns 1-4) shows results with Determinantal Point Process (DPP)-based inducing point selection [30], still in the case of asymmetric outliers.
+The bottom row (columns 1–4) shows results with Determinantal Point Process (DPP)-based inducing point selection [moss2023inducing], still in the case of asymmetric outliers.
 RCaGP remains the top performer on DNA and performs similarly to CaGP on other tasks, with RCSVGP slightly ahead on Hartmann.
-Substituting EI with Thompson Sampling in DPP-BO leads to equal performance across baselines (Figure S1).
-Next, Figure S2 presents results with TurBO, a BO variant tailored to high dimensions [8].
+Substituting EI with Thompson Sampling in DPP-BO leads to equal performance across baselines (Figure˜S1).
+Next, Figure˜S2 presents results with TurBO, a BO variant tailored to high dimensions [eriksson2019scalable].
 RCaGP exhibits strong average performance, and notably, it is the only method that never ranks last in a statistically significant manner.
@@
 Together, these results highlight RCaGP’s versatility and robustness across inducing point allocation strategies (EULBO, DPP), acquisition functions (EI, TS), BO behavior (global BO, trust-region BO), and noise regimes (with or without outliers).
-Its consistent superiority supports the central claim: combining robustness from RCGP with the uncertainty calibration of CaGP leads to a synergistic improvement—not just an additive one—across a wide range of BO challenges.
+Its consistent superiority supports the central claim: combining robustness from RCGP with the uncertainty calibration of CaGP leads to a synergistic improvement, not just an additive one, across a wide range of BO challenges.
 Figure 4: High-throughput Bayesian Optimization task.
@@
 See Appendix G for further analysis.
-8 Discussion
+8 Conclusion
 This paper introduced Robust Computation-Aware Gaussian Processes (RCaGP), a principled framework unifying robust inference and approximation-aware modeling coherently rather than tackling these challenges independently.
@@
 Lastly, broader evaluations across more complex domains remain an avenue for future work.
+Acknowledgments.
+This research was supported by the Research Council of Finland (flagship programme: Finnish Center for Artificial Intelligence, FCAI grants 358958, 345604, and 341763), and the UKRI Turing AI World-Leading Researcher Fellowship, EP/W002973/1.
+We also acknowledge the computational resources provided by the Aalto Science-IT Project.
 Supplementary Material - Robust and Computation-Aware Gaussian Processes
 The appendix is organized as follows:
-  • Appendix A gives the results of additional experiments mentioned in the main text.
+  • Appendix˜A gives the results of additional experiments mentioned in the main text.
 These include experiments on UCI regression datasets with uniform and focused outliers (Table S1 and Table S2), BO results for DPP-BO-TS (Figure S1), TuRBO (Figure S2), and dataset-specific results for the ablation study conducted in Section 7.4, where $c$ is varied.
-  • Appendix B gives the computational complexity of RCaGP.
-  • Appendix C formally derives the ELBO for RCaGP introduced in Section 3.2.
-  • Appendix D contains the proof for the robustness property presented in Proposition 4.1.
-  • Appendix E proves Proposition 4.2, showing that RCaGP’s uncertainty estimates capture the worse-case error over all latent functions.
-  • Appendix F contains additional theoretical results: a proposition showing the convergence in mean of RCaGP in RKHS norm (Section F.2), and a proposition showing the lack of robustness of CaGP (Section F.1).
-  • Appendix G provides further details regarding the expert-driven prior mean introduced in Section 5 and employed during the ablation study in Section 7.3.
-  • Appendix H provides additional details about the UCI regression datasets (Section H.1), the test functions employed in BO (Section H.2), and the outlier-contamination protocols studied in this work (Section H.3).
-  • Appendix I provides tables that report the hyperparmeters used for UCI regression and high-throughput BO experiments (Section I.1) and the description of the computing resources used for our experiments (Section I.2).
+  • Appendix˜B gives the computational complexity of RCaGP.
+  • Appendix˜C formally derives the ELBO for RCaGP introduced in Section 3.2.
+  • Appendix˜D contains the proof for the robustness property presented in Proposition˜4.1.
+  • Appendix˜E proves Proposition˜4.2, showing that RCaGP’s uncertainty estimates capture the worse-case error over all latent functions.
+  • Appendix˜F contains additional theoretical results: a proposition showing the convergence in mean of RCaGP in RKHS norm (Section F.2), and a proposition showing the lack of robustness of CaGP (Section F.1).
+  • Appendix˜G provides further details regarding the expert-driven prior mean introduced in Section 5 and employed during the ablation study in Section 7.4.
+  • Appendix˜H provides additional details about the UCI regression datasets (Section H.1), the test functions employed in BO (Section H.2), and the outlier-contamination protocols studied in this work (Section H.3).
+  • Appendix˜I provides tables that report the hyperparmeters used for UCI regression and high-throughput BO experiments (Section I.2) and the description of the computing resources used for our experiments (Section I.1).
 Appendix A Additional experiments
-The results of the UCI regression experiments in the presence of uniform and focused outliers are presented in Table S1 and Table S2, respectively.
+The results of the UCI regression experiments in the presence of uniform and focused outliers are presented in Table˜S1 and Table˜S2, respectively.
 The MAE and NLL metrics indicate that RCaGP consistently outperforms the baseline methods across most datasets.
 These findings demonstrate the robustness and versatility of RCaGP in handling different types of outliers.
+We further conducted experiments using the radial basis function (RBF) kernel on both clean and contaminated datasets.
+While RCaGP achieves competitive MAE and NLL performance compared to RCSVGP on clean data, it substantially outperforms all baselines on contaminated data.
 However, this performance improvement comes at the cost of increased computational time compared to the baselines.
-The results of Bayesian optimization (BO) experiments using the Thompson Sampling (TS) acquisition function (AF) and Trust-Region BO (TuRBO) are presented in Figure S1 and Figure S2, respectively.
+The results of Bayesian optimization (BO) experiments using the Thompson Sampling (TS) acquisition function (AF) and Trust-Region BO (TuRBO) are presented in Figure˜S1 and Figure˜S2, respectively.
 In experiments utilizing the TF acquisition function, no model demonstrates a statistically significant performance advantage.
@@
 This suggests that the presence of outliers in the Lunar task may impede TuRBO’s ability to effectively explore and identify the global optimum.
-The MAE and NLL plots for varying values of $\epsilon$ in the weight function $w$ are shown in Figure S3.
+The MAE and NLL plots for varying values of $\epsilon$ in the weight function $w$ are shown in Figure˜S3.
 Across all datasets, RCaGP consistently yields lower MAE and NLL values than RCSVGP, indicating superior predictive performance.
@@
 Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
+Bolded results refer to the best baseline.
+Lower is better.
 Boston Energy Yacht Parkinsons
@@
 Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
+Bolded results refer to the best baseline.
+Lower is better.
 Boston Energy Yacht Parkinsons
@@
 clock-time 2.570 ± 0.050 3.780 ± 0.110 0.930 ± 0.060 1.010 ± 0.040
-Figure S1: High-throughput Bayesian Optimization task under asymmetric outliers using DPP-BO and the Thompson Sampling (TS) acquisition strategy.
-Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std.
-Figure S2: High-throughput Bayesian Optimization task using TurBO under asymmetric outliers.
-Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std.
-Top row: EULBO-TuRBO-EI.
-Bottom row: DPP-TuRBO-EI.
-Figure S3: Dataset-specific results for the ablation study conducted in Section 7.3, varying $c$ in weight function $w$ (Equation 12).
+Table S3: UCI Regression datasets results with RBF kernel.
+Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
+Bolded results refer to the best baseline.
+Lower is better.
+Asymmetric outliers No outliers
+Boston Energy Yacht Parkinsons Boston Energy Yacht Parkinsons
+SVGP MAE 1.145 ± 0.052 1.138 ± 0.040 1.235 ± 0.044 1.002 ± 0.100 0.731 ± 0.052 0.908 ± 0.026 0.767 ± 0.063 0.811 ± 0.078
+NLL 1.691 ± 0.043 1.736 ± 0.037 1.732 ± 0.041 1.590 ± 0.079 1.433 ± 0.053 1.430 ± 0.023 1.420 ± 0.078 1.448 ± 0.070
+clock-time 0.930 ± 0.280 1.170 ± 0.070 0.650 ± 0.060 0.720 ± 0.260 0.940 ± 0.030 1.230 ± 0.150 0.870 ± 0.240 0.670 ± 0.130
+CaGP MAE 1.060 ± 0.095 0.971 ± 0.090 1.133 ± 0.092 0.955 ± 0.110 0.488 ± 0.042 0.334 ± 0.024 0.512 ± 0.042 0.674 ± 0.077
+NLL 1.611 ± 0.086 1.535 ± 0.071 1.679 ± 0.088 1.556 ± 0.080 1.278 ± 0.040 1.106 ± 0.011 1.263 ± 0.042 1.362 ± 0.067
+clock-time 2.470 ± 0.060 3.540 ± 0.080 0.640 ± 0.100 0.870 ± 0.080 2.410 ± 0.060 3.540 ± 0.160 0.770 ± 0.060 0.830 ± 0.100
+RCSVGP MAE 0.680 ± 0.122 0.456 ± 0.041 0.733 ± 0.114 0.840 ± 0.187 0.512 ± 0.074 0.558 ± 0.035 0.640 ± 0.074 0.540 ± 0.083
+NLL 1.364 ± 0.080 1.200 ± 0.026 1.381 ± 0.084 1.518 ± 0.116 1.302 ± 0.063 1.300 ± 0.034 1.318 ± 0.047 1.417 ± 0.065
+clock-time 1.500 ± 0.090 2.560 ± 0.100 0.970 ± 0.090 0.970 ± 0.090 1.530 ± 0.040 2.780 ± 0.120 1.130 ± 0.080 0.720 ± 0.060
+RCaGP MAE 0.733 ± 0.156 0.353 ± 0.053 0.651 ± 0.095 0.913 ± 0.198 0.379 ± 0.050 0.326 ± 0.023 0.365 ± 0.035 0.534 ± 0.078
+NLL 1.405 ± 0.118 1.173 ± 0.024 1.336 ± 0.061 1.531 ± 0.152 1.241 ± 0.046 1.144 ± 0.007 1.226 ± 0.034 1.307 ± 0.068
+clock-time 2.550 ± 0.060 3.590 ± 0.090 0.790 ± 0.060 0.900 ± 0.030 2.470 ± 0.030 3.740 ± 0.170 0.940 ± 0.070 1.070 ± 0.240
+Student-t MAE 0.962 ± 0.037 1.015 ± 0.028 0.907 ± 0.061 1.025 ± 0.079 0.896 ± 0.041 0.970 ± 0.025 0.815 ± 0.071 0.965 ± 0.074
+NLL 1.619 ± 0.039 1.606 ± 0.024 1.564 ± 0.067 1.658 ± 0.085 1.593 ± 0.063 1.572 ± 0.026 1.559 ± 0.100 1.627 ± 0.092
+clock-time 1.030 ± 0.040 1.350 ± 0.070 0.790 ± 0.060 0.730 ± 0.040 1.090 ± 0.030 1.460 ± 0.060 0.920 ± 0.050 0.730 ± 0.010
 Appendix B Computational Complexity
@@
 $\ell_{\mathrm{ELBO}}^{\mathrm{RCaGP}}={\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]-{\mathrm{KL}}[q_{i}({\bf{f}})\|p({\bf{f}})]$ (S3)
-$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}}))+\hat{\mu}_{i}({\bf{X}})^{\top}\nu$
+$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}}))+\hat{\mu}_{i}({\bf{X}})^{\top}\nu\$
 $\quad-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})-\frac{1}{2}(\bar{{\bf{v}}}_{i}^{\top}{\bf{S}}_{i}^{\top}{\bf{K}}{\bf{S}}_{i}\bar{{\bf{v}}}_{i}+\log\det({\bf{S}}_{i}^{\top}\tilde{{\bf{K}}}{\bf{S}}_{i})-i\log({\sigma_{\mathrm{noise}}^{2}})$
@@
 We first compute the expected loss function term:
-[…]
+${\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]$ $=\int\log\exp(-nL^{w}_{n}({\bf{f}},{\bf{y}},{\bf{X}}))q_{i}({\bf{f}})d{\bf{f}}$ (S6)
 $=\int-\frac{1}{2}({\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}-2{\bf{f}}^{\top}\nu+C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}}))q_{i}({\bf{f}})d{\bf{f}}$ (S7)
-[…]
-[…]
-[…]
-Next, we apply the identity ${\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}={\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})$ (see [2]):
+$=\int-\frac{1}{2}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}\,q_{i}({\bf{f}})\,d{\bf{f}}+\int{\bf{f}}^{\top}\nu\,q_{i}({\bf{f}})\,d{\bf{f}}$
+$\quad-\int\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})\,q_{i}({\bf{f}})d{\bf{f}}$ (S8)
+$=-\frac{1}{2}\int{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}\,q_{i}({\bf{f}})\,d{\bf{f}}+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}^{\top}\nu]-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}}),$ (S9)
+Next, we apply the identity ${\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}={\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})$ (see [altamirano2023robust]):
 ${\mathbb{E}}_{q_{i}({\bf{f}})}[\log p^{w}({\bf{y}}|{\bf{f}})]$ $=-\frac{1}{2}{\mathbb{E}}_{q_{i}({\bf{f}})}[{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})]+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}^{\top}\nu]-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})$ (S10)
@@
 $(\hat{{\bf{S}}}\hat{{\bf{B}}}\hat{{\bf{S}}}^{\top})^{-1}$ $=\hat{{\bf{S}}}^{+\top}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{G}}}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{+}$ (S28)
-where we define $\hat{{\bf{G}}}={\bf{I}}-\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}})(\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}}))^{+}$ and + denotes the Moore-Penrose inverse.
+where we define $)+\hat{{\bf{G}}}={\bf{I}}-\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}})(\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}}$$)^{+}$ and + denotes the Moore-Penrose inverse.
 Proof:
@@
 $(\hat{{\bf{S}}}\hat{{\bf{B}}}\hat{{\bf{S}}}^{\top})^{-1}=\hat{{\bf{S}}}^{+\top}\hat{{\bf{B}}}^{-1/2}\underbrace{({\bf{I}}-\hat{{\bf{B}}}^{-1/2}\hat{{\bf{O}}}(\hat{{\bf{B}}}^{-1/2}\hat{{\bf{O}}})^{+})}_{\hat{{\bf{G}}}}\hat{{\bf{B}}}^{-1/2}\hat{{\bf{S}}}^{+}$ (S38)
-PIF of RCaGP We now prove Proposition 4.1: Let ${\bf{f}}\sim{\mathcal{GP}}(m,k)$ denote the RCaGP prior, and let $i\in{0,\dots,n}$ represent the number of actions in RCaGP.
+PIF of RCaGP We now prove Proposition˜4.1: Let ${\bf{f}}\sim{\mathcal{GP}}(m,k)$ denote the RCaGP prior, and let $i\in{0,\dots,n}$ represent the number of actions in RCaGP.
 Assume the observation noise is $\varepsilon\sim{\mathcal{N}}(\mathbf{0},{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})$.
@@
 $=\left(\sum_{j=1}^{n}\left(4\sigma_{\mathrm{noise}}^{-2}w^{2}({\bf{x}}_{j},y_{j})\right)\sum_{k=1}^{n}\left({\bf{K}}_{kk}+\sigma^{4}_{\mathrm{noise}}/2\;w^{-2}({\bf{x}}_{k},y_{k})\right)\right)\bar{C}_{1}\bar{C}_{2}-n$ (S60)
-[…]
-Next, we derive the bound for the term (2). Following [2], we have that
+$\leq\left(n\sup_{{\bf{x}},y}\,4\sigma^{-2}_{\mathrm{noise}}\,w^{2}({\bf{x}},y)\,\left(n\sup_{\hat{{\bf{x}}},\hat{y}}\sigma^{4}_{\mathrm{noise}}/2\;w^{-2}(\hat{{\bf{x}}},\hat{y})+\sum_{k=1}^{n}{\bf{K}}_{kk}\right)\right)\bar{C}_{1}\bar{C}_{2}-n=\bar{C}_{3}$ (S61)
+Next, we derive the bound for the term (2). Following [altamirano2023robust], we have that
 $(2)\leq\lambda_{\max}((\hat{{\bf{K}}}_{i}^{c})^{-1})\|\hat{{\boldsymbol{\mu}}}_{i}^{c}-\hat{{\boldsymbol{\mu}}}_{i}\|_{1}^{2},$ (S62)
@@
 Finally, the last inequality holds since $\tilde{{\bf{K}}}_{i}=(\tilde{{\bf{C}}}_{i}^{-1}-\tilde{{\boldsymbol{\Sigma}}}_{i})$.
-Applying results from [2], we obtain
+Applying results from [altamirano2023robust], we obtain
 $(2)$ $\leq\,q\,\bar{C}_{6}\,\|{\bf{K}}\|_{F}\,\|({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{{\bf{w}}^{c}})^{-1}({\bf{y}}-{\bf{m}}_{{\bf{w}}^{c}})-({\bf{K}}+{\sigma_{\mathrm{noise}}^{2}}{\bf{J}}_{\bf{w}})({\bf{y}}-{\bf{m}}_{\bf{w}})\|_{1}^{2}$ (S75)
-[…]
+$\leq\,q\,\bar{C}_{6}\,\|{\bf{K}}\|_{F}\,2((\bar{C}_{7}+\bar{C}_{8})^{2}+(\bar{C}_{9}+\bar{C}_{10})^{2}(w(x_{n},y_{n}^{c})^{2}y_{n}^{c})^{2})$ (S76)
 $\leq\,\bar{C}_{11}+\bar{C}_{12}(w(x_{n},y_{n}^{c})^{2}y_{n}^{c})^{2}$ (S77)
 where $\bar{C}_{11}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{7}+\bar{C}_{8})^{2}$ and $\bar{C}_{12}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{9}+\bar{C}_{10})^{2}$.
-The terms $\bar{C}_{7},\bar{C}_{8},\bar{C}_{9},\bar{C}_{10}$ correspond to $\tilde{C}_{6},\tilde{C}_{8},\tilde{C}_{7},\tilde{C}_{9}$ in [2].
+The terms $\bar{C}_{7},\bar{C}_{8},\bar{C}_{9},\bar{C}_{10}$ correspond to $\tilde{C}_{6},\tilde{C}_{8},\tilde{C}_{7},\tilde{C}_{9}$ in [altamirano2023robust].
 The term (3) can be written as follows:
@@
 $\delta_{w}({\bf{x}},{\bf{x}}^{\prime})=\begin{array}[]{cc}&\begin{cases}w^{-2}({\bf{x}},y)&{\bf{x}}={\bf{x}}^{\prime}\,\text{and}\,{\bf{x}}\in{\mathcal{D}}\\ 2&{\bf{x}}={\bf{x}}^{\prime}\,\text{and}\,{\bf{x}}\notin{\mathcal{D}}\\ 0&{\bf{x}}\neq{\bf{x}}^{\prime}\end{cases}\end{array},$ (S90)
-where ${\mathcal{D}}$ denotes the observation set. Then, applying [20, Lemma 3.9] provides
+where $\mathcal{D}$} denotes the observation set. Then, applying [kanagawa2018gaussian, Lemma 3.9] provides
 $\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}h({\bf{x}})-\hat{\mu}_{i}^{g}({\bf{x}})\right)^{2}=\left(\sup_{\|h\|_{{\mathcal{H}}_{k^{w}}}\leq 1}h({\bf{x}})-\sum_{j=1}^{n}c_{j}h({\bf{x}}_{j})\right)^{2}$ (S91)
@@
 since ${\bf{x}}\neq{\bf{x}}_{j}$, it holds that $k^{w}({\bf{x}},{\bf{X}})=k({\bf{x}},{\bf{X}})$.
-By definition, we have $k^{w}({\bf{X}},{\bf{X}})=\tilde{{\bf{K}}}={\bf{K}}+\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}{\bf{J}}_{\bf{w}}$, and following in [45, Eq.
+By definition, we have $k^{w}({\bf{X}},{\bf{X}})=\tilde{{\bf{K}}}={\bf{K}}+\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}{\bf{J}}_{\bf{w}}$, and following in [wenger2022posterior, Eq.
 (S42)], it holds that $\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}=\tilde{{\bf{C}}_{i}}$.
@@
 $=k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{K}}}^{-1}\tilde{{\bf{K}}}\tilde{{\bf{K}}}^{-1}k^{w}({\bf{X}},{\bf{x}})-2k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{K}}}^{-1}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})+$
-$\quad\;k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})$ (S100)
+$\qquad\;k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})$ (S100)
 $=k({\bf{x}},{\bf{X}})(\tilde{{\bf{K}}}^{-1}-\tilde{{\bf{C}}}_{i})k({\bf{X}},{\bf{x}})$ (S101)
@@
 $\frac{\hat{\mu}({\bf{x}})-\hat{\mu}_{i}({\bf{x}})}{\|h\|_{{\mathcal{H}}_{k^{w}}}}\leq\sqrt{\tilde{\sigma}^{2}_{i}({\bf{x}},{\bf{x}})}$ (S104)
-The proof follows immediately from Proposition 4.2 and notice that $h/\|h\|_{{\mathcal{H}}_{k^{w}}}$ has unit norm.
-Corollary E.1 implies that the computational uncertainty provides the pointwise bound relative to RCGP’s mathematical posterior, and the combined uncertainty is the pointwise bound relative to the shifted latent function.
+The proof follows immediately from Proposition˜4.2 and notice that $h/\|h\|_{{\mathcal{H}}_{k^{w}}}$ has unit norm.
+Corollary˜E.1 implies that the computational uncertainty provides the pointwise bound relative to RCGP’s mathematical posterior, and the combined uncertainty is the pointwise bound relative to the shifted latent function.
 Appendix F Additional Theoretical Results
@@
 $\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)=\frac{1}{2}\left(\mathrm{Tr}({\bf{K}}_{i}^{c}{\bf{K}}_{i})-n+({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})^{\top}({\bf{K}}_{i}^{c})^{-1}({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})+\ln\left(\frac{\mathrm{det}({\bf{K}}_{i}^{c})}{\mathrm{det}({\bf{K}}_{i})}\right)\right)$ (S110)
-Based on [2], the PIF leads to the following form:
+Based on [altamirano2023robust], the PIF leads to the following form:
 $\mathrm{PIF}_{\mathrm{CaGP}}(y_{m}^{c},{\mathcal{D}},i)=\frac{1}{2}\left(({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})^{\top}({\bf{K}}_{i}^{c})^{-1}({\boldsymbol{\mu}}_{i}^{c}-{\boldsymbol{\mu}}_{i})\right)$ (S111)
@@
 We first show the corresponding empirical-risk minimization problem of RCGP whenever ${\bf{m}}=\mathbf{0}$.
-Following [2, proof of Proposition 3.1], we can rewrite $L_{n}^{w}$ and formulate the RCGP objective as follows:
+Following [altamirano2023robust, proof of Proposition 3.1], we can rewrite $L_{n}^{w}$ and formulate the RCGP objective as follows:
 $\hat{{\bf{f}}}=\mathrm{argmin}_{{\bf{f}}\in{\mathcal{H}}_{k}}\frac{1}{2n}\left(\underbrace{{\bf{f}}^{\top}\lambda^{-1}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}-2{\bf{f}}^{\top}\lambda^{-1}{\bf{J}}_{\bf{w}}^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})}_{L^{w}_{n}}\right)+\frac{1}{2}\|{\bf{f}}\|^{2}_{{\mathcal{H}}_{k}}$ (S119)
@@
 Proof:
-The optimization problem in S119 allows us to apply the representer theorem [36]. It implies that the solution of S119 can be written as a weighted sum, i.e.,
+The optimization problem in S119 allows us to apply the representer theorem [scholkopf2001generalized]. It implies that the solution of S119 can be written as a weighted sum, i.e.,
 $\hat{{\bf{f}}}=\sum_{j=1}^{n}\hat{\alpha}_{j}k(.,{\bf{x}}_{j})$ (S122)
@@
 Therefore, we conclude that $\hat{{\boldsymbol{\alpha}}}=({\bf{K}}+\lambda{\bf{J}}_{\bf{w}})^{-1}({\bf{y}}-{\bf{m}}_{\bf{w}})$ provides the unique solution to S119.
-As a remark, Proposition F.1 closely connects with [20, Theorem 3.4].
+As a remark, Proposition F.1 closely connects with [kanagawa2018gaussian, Theorem 3.4].
 F.2.2 Proof of Proposition F.3
-Relative bound errors. We first provide the equivalence of [45, Proposition 2]:
+Relative bound errors. We first provide the equivalence of [wenger2022posterior, Proposition 2]:
 Proposition F.2.
@@
 where $\bar{{\bf{v}}}=\hat{{\bf{v}}}/\|\tilde{{\bf{v}}}\|_{\tilde{{\bf{K}}}}$.
-The proof is by directly substituting ${\bf{C}}_{i},\hat{{\bf{K}}},{\bf{v}}_{\ast}$ in [45] with $\tilde{{\bf{C}}}_{i},\tilde{{\bf{K}}},\hat{{\bf{v}}}$, respectively.
+The proof is by directly substituting ${\bf{C}}_{i},\hat{{\bf{K}}},{\bf{v}}_{\ast}$ in [wenger2022posterior] with $\tilde{{\bf{C}}}_{i},\tilde{{\bf{K}}},\hat{{\bf{v}}}$, respectively.
 Next, we establish the convergence theorem of RCaGP through the following proposition:
@@
 The third inequality stems from the definition of ${\bf{J}}_{\bf{w}}$ and the fact that the maximum eigenvalue of a diagonal matrix is the largest component of its diagonal.
-Assuming $\hat{\mu}_{i}(.)=m(.)+\sum_{j=1}^{n}(\tilde{{\bf{v}}}_{i})_{j}k(.,{\bf{x}}_{j})=m(.)+k(.,{\bf{X}})\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and applying result from [45], we have that
+Assuming $\hat{\mu}_{i}(.)=m(.)+\sum_{j=1}^{n}(\tilde{{\bf{v}}}_{i})_{j}k(.,{\bf{x}}_{j})=m(.)+k(.,{\bf{X}})\tilde{{\bf{C}}}_{i}({\bf{y}}-{\bf{m}}_{\bf{w}})$ and applying result from [wenger2022posterior], we have that
 $\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|_{\bf{K}}^{2}=\|\hat{{\boldsymbol{\mu}}}_{\ast}-\hat{{\boldsymbol{\mu}}}_{i}\|^{2}_{{\mathcal{H}}_{k}}$ (S132)
@@
 $\|\hat{{\boldsymbol{\mu}}}_{\ast}-\hat{{\boldsymbol{\mu}}}_{i}\|_{{\mathcal{H}}_{k}}=\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{i}\|_{\bf{K}}\leq\hat{\rho}(i)c({\bf{J}}_{\bf{w}})\|\hat{{\bf{v}}}-\tilde{{\bf{v}}}_{0}\|_{\bf{K}}=\hat{\rho}(i)c({\bf{J}}_{\bf{w}})\|\hat{{\boldsymbol{\mu}}}_{\ast}-{\bf{m}}\|_{{\mathcal{H}}_{k}}$ (S133)
+F.3 SVGP with Relevance Pursuit
+Proposition F.4.
+Let ${\mathcal{D}}_{\setminus i}=\{({\bf{x}}_{j},y_{j}):j\neq i\},{\bf{Z}}={z_{m}}$ is the inducing points, ${\boldsymbol{\rho}}={\boldsymbol{\rho}}_{\setminus i}+\rho_{i}\mathbf{e}_{i}$, where ${\boldsymbol{\rho}},{\boldsymbol{\rho}}_{\setminus i}\in_{+}^{n}$, ${\boldsymbol{\rho}}_{\setminus i}]_{i}=0$, and $\mathbf{e}_{i}$ is the ith canonical basis vector.
+Then keeping ${\boldsymbol{\rho}}_{\setminus i}$ fixed,
+$\rho_{i}^{\ast}={\mathrm{argmax}}_{\rho_{i}}\,\mathrm{ELBO}({\boldsymbol{\rho}}_{\setminus i}+\rho_{i}\mathbf{e}_{i})=\left[\frac{1}{\sqrt{{\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}}}(y_{i}-{\mathbb{E}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}])-{\mathbb{V}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}]\right]_{+}$ (S134)
+where $y_{i}=f({\bf{x}}_{i})+\epsilon_{i}$. The quantities can be expressed as a function of ${\bf{A}}^{-1}=({\bf{Q}}_{\bf{f}}+{\bf{D}}_{{\sigma_{\mathrm{noise}}^{2}}+{\boldsymbol{\rho}}})^{-1}$ :
+${\mathbb{E}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}]=y_{i}-[{\bf{A}}^{-1}{\bf{y}}]_{i}/[{\bf{A}}^{-1}]_{ii}\quad\text{and}\quad{\mathbb{V}}_{q}[y_{i}|{\mathcal{D}}_{\setminus i}]=1/[{\bf{A}}^{-1}]_{ii}$ (S135)
+where ${\bf{D}}_{{\sigma_{\mathrm{noise}}^{2}}+{\boldsymbol{\rho}}}$ is a diagonal matrix whose entries are ${\sigma_{\mathrm{noise}}^{2}}+{\boldsymbol{\rho}}$.
+Proof: Following [titsias2009variational], the evidence lower bound (ELBO) of GP can be written as
+$2\mathrm{ELBO}({\boldsymbol{\theta}})=n\log(2\pi)+\log\det(\mathbf{A})+{\bf{y}}^{\top}{\bf{A}}^{-1}{\bf{y}}+{\mathrm{tr}}({\bf{D}}_{{\boldsymbol{\rho}}+{\sigma_{\mathrm{noise}}^{2}}}^{-1}({\bf{K}}-{\bf{Q}}_{\bf{f}})),$ (S136)
+where ${\bf{A}}={\bf{Q}}_{\bf{f}}+{\bf{D}}_{{\boldsymbol{\rho}}+{\sigma_{\mathrm{noise}}^{2}}}$, and ${\bf{Q}}_{\bf{f}}={\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}$
+Following [ament2024robust], we partition matrix $\bf{A}$} to separate the effect of $\rho_{i}$ and use Schur’s complement:
+${\bf{A}}^{-1}=\begin{bmatrix}{\bf{A}}^{-1}_{\setminus i}+\hat{{\bf{u}}}\beta_{i}\hat{{\bf{u}}}^{\top}&-\hat{{\bf{u}}}\beta_{i}\\ -\hat{{\bf{u}}}^{\top}\beta_{i}&\beta_{i}\end{bmatrix}$, (S137)
+where
+${\bf{A}}_{\setminus i}=[{\bf{Q}}_{\bf{f}}]_{\setminus i}+{\bf{D}}_{{\boldsymbol{\rho}}_{\setminus i}+{\sigma_{\mathrm{noise}}^{2}}},$ (S138)
+$\hat{{\bf{u}}}={\bf{A}}^{-1}_{\setminus i}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}),$ (S139)
+$\hat{\beta}_{i}=([k({\bf{x}}_{i},{\bf{x}}_{i})+{\sigma_{\mathrm{noise}}^{2}}+\rho_{i}]-k({\bf{x}}_{i},{\bf{Z}}){\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}{\bf{A}}_{\setminus i}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}))^{-1}.$ (S140)
+Quadratic term:
+${\bf{y}}^{\top}({\bf{A}}+{\bf{D}}_{{\boldsymbol{\rho}}+{\sigma_{\mathrm{noise}}^{2}}})^{-1}{\bf{y}}={\bf{y}}^{\top}_{\setminus i}{\bf{A}}^{-1}_{\setminus i}{\bf{y}}_{\setminus i}+\beta_{i}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}.$ (S141)
+Determinant term:
+$\det({\bf{A}})=\det({\bf{Q}}_{\bf{f}}+{\bf{D}}_{{\boldsymbol{\rho}}_{i}+{\sigma_{\mathrm{noise}}^{2}}})=\det({\bf{A}}_{\setminus i})\beta_{i}^{-1}.$ (S142)
+Following [ament2024robust], we formulate the ELBO difference as
+$2(\mathrm{ELBO}({\boldsymbol{\rho}})-\mathrm{ELBO}({\boldsymbol{\rho}}_{\setminus i}))=-\hat{\beta}_{i}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}-\log(\hat{\beta}_{i}^{-1})-\rho_{i}({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}).$ (S143)
+The derivative of the difference in ELBO w.r.t. $\rho_{i}$, is
+$\partial_{\rho_{i}}2(\mathrm{ELBO}({\boldsymbol{\rho}})-\mathrm{ELBO}({\boldsymbol{\rho}}_{\setminus i}))=({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}\hat{\beta}_{i}^{2}-\hat{\beta}_{i}-({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}).$ (S144)
+Using the quadratic formula, we obtain the solution $\hat{\beta}_{1,2}=\frac{1\pm\sqrt{1+4({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})}}{2({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}}$.
+Note that $({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})^{2}>0$ and ${\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii}\geq 0$.
+Therefore, the smaller root $\rho_{2}$ is negative when $({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})>0$ and zero when $({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})=0$.
+Based on the $\hat{\beta}$ formulation (predictive variance cannot be zero), we reject $\hat{\beta}_{2}<0$.
+Similarly, we reject $\hat{\beta}_{2}=0$, following [ament2024robust].
+By ignoring the constant term in $\hat{\beta}_{2}$ and following the approach of [ament2024robust], we solve $\hat{\beta}_{2}^{-1}=1/\sqrt{({\bf{K}}_{ii}-[{\bf{Q}}_{\bf{f}}]_{ii})}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})$ for $\rho_{i}$ and projecting to non-negative half-line, we get
+$\rho_{i}=[\frac{1}{\sqrt{({\bf{K}}_{ii}}-[{\bf{Q}}_{\bf{f}}]_{ii})}({\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i})-(k({\bf{x}}_{i},{\bf{x}}_{i})+{\sigma_{\mathrm{noise}}^{2}}-k({\bf{x}}_{i},{\bf{Z}}){\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}{\bf{A}}_{\setminus i}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}))]_{+}.$ (S145)
+Lastly, we note that
+${\bf{y}}_{\setminus i}^{\top}\hat{{\bf{u}}}-y_{i}={\bf{y}}_{\setminus i}^{\top}({\bf{A}}^{-1}_{\setminus i}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}}))-y_{i}={\mathbb{E}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}]-y_{i},$ (S146)
+$k({\bf{x}}_{i},{\bf{x}}_{i})+{\sigma_{\mathrm{noise}}^{2}}-k({\bf{x}}_{i},{\bf{Z}}){\bf{K}}_{\bf{z}}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}^{\top}{\bf{A}}_{\setminus i}^{-1}{\bf{K}}_{{\bf{x}}{\bf{z}}}{\bf{K}}_{\bf{z}}^{-1}k^{\top}({\bf{x}}_{i},{\bf{Z}})={\mathbb{V}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}].$ (S147)
+Following [Rasmussen2006], we have that
+${\mathbb{E}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}]=y_{i}-[{\bf{A}}^{-1}{\bf{y}}]_{i}/[{\bf{A}}^{-1}]_{i},$ (S148)
+${\mathbb{V}}_{q}[{\bf{y}}_{i}|{\mathcal{D}}_{\setminus i}]=1/[{\bf{A}}^{-1}]_{ii},$ (S149)
+which follow the LOO predictive values.
 Appendix G Details of Expert-guided Robust Mean Prior
@@
 For every outlier candidate $({\bf{x}}_{o},\hat{y}_{o})$, we have that
-[…]
-[…]
+$p(\bar{\delta}_{o})=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})$ $\text{prior},$ (S150)
+$p(\bar{o}_{o}\,|\,\bar{\delta}_{o})=\bar{\delta}_{o}\bar{o}_{o}+(1-\bar{\delta}_{o})(1-\bar{o}_{o})$ $\text{likelihood},$ (S151)
 $p(\bar{\delta}_{o}|\bar{o}_{o})\propto p(\bar{\delta}_{o})\,p(\bar{o}_{o}|\bar{\delta}_{o})$ posterior
-$\qquad\quad\;\;=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})\left(\bar{\delta}_{o}(\bar{o}_{o})+(1-\bar{\delta}_{o})(1-\bar{o}_{o})\right)$
-$\qquad\quad\;\;=\mathcal{B}(\alpha=\alpha_{o}+\bar{o}_{o},\beta=\beta_{o}+1-\bar{o}_{o}),$ (S136)
+$\hskip 17.00024pt\qquad\;\;=\mathcal{B}(\alpha=\alpha_{o},\beta=\beta_{o})\left(\bar{\delta}_{o}(\bar{o}_{o})+(1-\bar{\delta}_{o})(1-\bar{o}_{o})\right)$
+$\hskip 17.00024pt\qquad\;\;=\mathcal{B}(\alpha=\alpha_{o}+\bar{o}_{o},\beta=\beta_{o}+1-\bar{o}_{o}),$ (S152)
 where we define the hyperparameters $\alpha_{0}$ and $\beta_{0}$ as follows:
@@
 Next, we outline the posterior inference details for the outlier corrections provided by the human expert:
-[…]
-[…]
+$p(\bar{\mu}_{o})={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})$ $\text{prior},$ (S154)
+$p(\bar{y}_{o}|\bar{\mu}_{o})={\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ $\text{likelihood},$ (S155)
 $p(\bar{\mu}_{o}|\bar{y}_{o})\propto p(\bar{\mu}_{o})\,p(\bar{y}_{o}|\bar{\mu}_{o})$ posterior (S156)
-$\qquad\quad\;\;={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})\,{\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ (S141)
-$\qquad\quad\;\;={\mathcal{N}}\left(\mu=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},\sigma^{2}=(\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}})^{-1}\right),$ (S142)
+$\hskip 17.00024pt\qquad\;\;={\mathcal{N}}(\mu=\mu_{o},\sigma^{2}=\tau_{o}^{-1})\,{\mathcal{N}}(\mu=\bar{\mu}_{o},\sigma^{2}=\sigma^{2}_{\mathrm{corr.}})$ (S157)
+$\hskip 17.00024pt\qquad\;\;={\mathcal{N}}\left(\mu=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},\sigma^{2}=(\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}})^{-1}\right),$ (S158)
 where we define the hyperparameters $\mu_{o}$ and $\tau_{o}$ as follows:
@@
 ${\mathbb{E}}_{p(\bar{\mu}_{o}|\bar{y}_{o})}=\frac{\tau_{o}\,\mu_{o}+\sigma^{-2}_{\mathrm{corr.}}\bar{y}_{o}}{\tau_{o}+1.\sigma^{-2}_{\mathrm{corr.}}},$ (S159)
-which equals the posterior mean. The algorithm to implement the informative mean prior is presented in Algorithm 1.
+which equals the posterior mean. The algorithm to implement the informative mean prior is presented in Algorithm˜1.
 Algorithm 1 Expert-guided robust mean prior algorithm
@@
 We try to predict the median price of homes residents own (excluding rented properties).
-The dataset can be found at https://www.cs.toronto.edu/~delve/data/boston/bostonDetail.html.
+The dataset can be found at https://www.cs.toronto.edu/˜delve/data/boston/bostonDetail.html.
 Energy
@@
 Hartmann 6D.
-The widely used Hartmann benchmark function [38].
+The widely used Hartmann benchmark function [surjanovic2020virtual].
 Lunar Lander.
@@
 The final objective value we optimize is the reward obtained by the policy averaged over a set of 50 random landing terrains.
-For this task, we use the same controller setup used by [8].
+For this task, we use the same controller setup used by [eriksson2019scalable].
 Rover.
-The rover trajectory optimization task introduced by [44] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a set of obstacles.
-We use the same obstacle set up as in [29].
+The rover trajectory optimization task introduced by [wang2018batched] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a set of obstacles.
+We use the same obstacle set up as in [robot].
 Lasso DNA.
-We optimize the $180-$dimensional DNA task from the LassoBench library [49] of benchmarks based on weighted LASSO regression.
+We optimize the $180$-dimensional DNA task from the LassoBench library [sehic2022lassobench] of benchmarks based on weighted LASSO regression.
 H.3 Additional outlier contamination protocols on UCI regression datasets
-Our outlier contamination protocols follow the same settings as [2]. We recall the asymmetric outlier case described in the main text in Section 7.
+Our outlier contamination protocols follow the same settings as [altamirano2023robust]. We recall the asymmetric outlier case described in the main text in Section 7.
 Asymmetric
@@
 Appendix I Experimental Details
-I.1 Hyperparameters
+I.1 Hardware
+For the UCI regression experiments, all models—including our proposed method and the baselines—were executed on a compute cluster consisting of two machines: one equipped with dual AMD EPYC 7713 processors (64 cores each, 2.0 GHz), and another with dual Intel Xeon Gold 6148 processors (20 cores each, 2.4 GHz).
+The longest-running UCI task, based on the Energy dataset, required approximately 15 minutes to complete, including both our model and all baseline methods.
+Bayesian Optimization experiments were conducted on a cluster with four NVIDIA V100 GPUs, each with 32 GB of memory.
+Among these tasks, the Hartmann 6D benchmark was the fastest, completing in roughly 20 minutes, while the most computationally intensive task—DNA—required up to 20 hours of runtime.
+I.2 Hyperparameters
 UCI regression
@@
 $\epsilon$ $0.2$
-$p$-outliers $0.1$
+pp-outliers $0.1$
 $n$-inducing points (SVGP) 100
@@
 $\sigma^{2}_{\mathrm{corr.}}$ $1.0$
-$p-$outliers 0.25
+pp-outliers 0.25
 Mean-prior $m$ (all models) $0.0$
@@
 Table S4: Hyperparameter settings used for RCaGP. Most of them remain consistent across all the tasks.
-I.2 Hardware
-For the UCI regression experiments, all models—including our proposed method and the baselines—were executed on a compute cluster consisting of two machines: one equipped with dual AMD EPYC 7713 processors (64 cores each, 2.0 GHz), and another with dual Intel Xeon Gold 6148 processors (20 cores each, 2.4 GHz).
-The longest-running UCI task, based on the Energy dataset, required approximately 15 minutes to complete, including both our model and all baseline methods.
-Bayesian Optimization experiments were conducted on a cluster with four NVIDIA V100 GPUs, each with 32 GB of memory.
-Among these tasks, the Hartmann 6D benchmark was the fastest, completing in roughly 20 minutes, while the most computationally intensive task—DNA—required up to 20 hours of runtime.
+Figure S1: High-throughput Bayesian Optimization task under asymmetric outliers using DPP-BO and the Thompson Sampling (TS) acquisition strategy.
+Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std.
+Figure S2: High-throughput Bayesian Optimization task using TurBO under asymmetric outliers.
+Each panel shows the best value found each iteration found so far, averaged across 20 repetitions $\pm$ 1 std.
+Top row: EULBO-TuRBO-EI.
+Bottom row: DPP-TuRBO-EI.
+Figure S3: Dataset-specific results for the ablation study conducted in Section 7.4, varying $c$ in weight function $w$ (Equation 12).
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The claims made in the abstract and introduction are theoretical, methodological, and empirical, and these are all backed in dedicated sections of the paper.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: A limitation paragraph has been added to the conclusion.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Yes, each proposition clearly states its theoretical assumptions. Proofs are deferred to the appendix.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Companion code along with running instructions will be provided in the supplementary materials.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: All benchmarks datasets employed for regression tasks are known, publicly available online benchmarks.
+URLs are provided in the Appendix.
+The test functions employed for Bayesian Optimization have been introduced in earlier works introduced in the Appendix, and will be available in the code we submit in Supplementary materials.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We provide the experimental details in Appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: We report the mean and standard deviation for every numerical experiment computed across a large number of train/test split in regression tasks, or across 20 repetitions with different seeds in Bayesian Optimization tasks.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We describe our computing resources in Appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have read all the regulations and make sure to follow them.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49.
+Justification: This work is foundational in nature and proposes a general-purpose Gaussian Process framework.
+It does not involve sensitive data, individual-level prediction, or deployment in decision-critical settings.
+As such, it does not pose direct societal risks or impacts.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [No]
+  54.
+Justification: Our work focuses on Gaussian Process methods and does not involve the release of models or datasets that pose a high risk of misuse.
+Therefore, safeguards for responsible release are not applicable.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite the original creators of the codebase that we use for the experiments in the paper.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [No]
+  64. Justification: We do not introduce any new asset.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [No]
+  69. Justification: We do not conduct research with human subjects or crowdsourcing.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [No]
+  74. Justification: Our research does not involve human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.